### v2

# Arcade Processes for Informed Martingale Interpolation

Abstract
Arcade processes are a class of continuous stochastic processes that interpolate in a strong sense, i.e., omega by omega, between zeros at fixed pre-specified times. Their additive randomization allows one to match any finite sequence of target random variables, indexed by the given fixed dates, on the whole probability space. The randomized arcade processes (RAPs) can thus be interpreted as a generalization of anticipative stochastic bridges. The filtrations generated by these processes are utilized to construct a class of martingales which interpolate between the given target random variables. These so-called filtered arcade martingales (FAMs) are almost-sure solutions to the martingale interpolation problem and reveal an underlying stochastic filtering structure. In the special case of conditionally-Markov randomized arcade processes, the dynamics of FAMs are informed by Bayesian updating. The same ideas are applied to filtered arcade reverse-martingales, which are constructed in a similar fashion, using reverse-filtrations of RAPs, instead. As a potential application of this theory, optimal transportation is explored: FAMs may be used to introduce noise in martingale optimal transport, in a similar fashion to how Schrödinger’s problem introduces noise in optimal transport. This information-based approach to transport is concerned with selecting an optimal martingale coupling for the target random variables under the influence of the noise that is generated by an arcade process, and suggests application in finance or climate science, for instance.
Keywords: Stochastic interpolation, stochastic bridges, martingale interpolation, martingale fitting, optimal transport, nonlinear stochastic filtering, information-based approach, Markov processes, Bayesian updating.
AMS2020: 49Q22, 60G35, 60G44, 60J60, 62M20, 94A15.
1 Introduction
The problem of constructing martingales which match given marginales has been the subject of many works in probability theory and applications thereof. The origin of the problem dates back to Strassen [29], with preliminary work by Blackwell [4], Sherman [26], and Stein [28], see [17] for a review. It was shown that a discrete-time martingale $(M_{n})_{n\in\mathbb{N}}$ can match a sequence of real probability measures $(\mu_{n})_{n\in\mathbb{N}}$ if and only if the measures $(\mu_{n})_{n\in\mathbb{N}}$ are convexly ordered, i.e., $M_{n}\sim\mu_{n}$ is possible for all $n\in\mathbb{N}$ if and only if $\int f(x)\,\mathrm{d}\mu_{n}(x)$ is an increasing sequence in $n$ for any positive convex function $f$. This result also extends to the continuous-time setting, see Kellerer’s theorem [16].
Strassen’s and Kellerer’s theorems do not provide a method for constructing the matching martingales. Many techniques are dedicated to Kellerer’s setting, where one is looking to construct a martingale that matches a given peacock [11]: Skorokhod’s embedding problem, Brownian random time-changes, local volatility models, scaling peacocks method, etc., see [21]. Since the target is a dynamical measure, the matching is produced in a weak sense, and the martingale evolves in continuous time. On the other hand, the matching in Strassen’s setting can be weak or almost sure, depending on the nature of the targets (measures or random variables), and the martingale is in discrete time. A widely known approach for this task is martingale optimal transport (MOT, [2], [3]). Inspired by Kantorovich’s optimal transport theory ([13], [31], [5]), it provides a framework for selecting a martingale coupling by minimizing a cost functional over the set of all martingale couplings.
Our interest lies at the intersection of Strassen’s and Kellerer’s settings: how to construct a continuous-time martingale that matches, in law or almost surely, a given finite set of convexly ordered random variables, at pre-specified dates? This is for instance of interest in a financial setting, where knowledge of the prices of vanilla call and put options provides an implied distribution for the underlying asset price at future fixed dates under a no-arbitrage requirement. This is an example of what is called the martingale interpolation problem, since one wishes to interpolate with a continuous-time martingale between the components of a discrete-time martingale. Imposing the martingale property on the interpolating process is the challenging part of this problem. If one put the martingale condition aside, then bridging the gap between a coupling and an interpolating measure is usually a manageable proposition. In Kantorovich’s optimal transport theory, for example, it is well-known that there is a one-to-one correspondence between optimal couplings and optimal interpolating measures, that is, geodesics in Wasserstein spaces, called McCann’s interpolations, under mild conditions. However, this property is not present in the martingale counterpart of optimal transport. Solutions to the martingale interpolation problem exist. The martingale Benamou-Brenier problem (MBBP) [1] aims to solve the optimization problem
$\sup\limits_{\mathcal{M}_{t}(\mu,\nu,B)}\mathbb{E}\left[\int_{0}^{1}\sigma_{t}\,\mathrm{d}t\right]$
over the set
$\mathcal{M}_{t}(\mu,\nu,B)$ =$=\left\{(\sigma_{t})_{t\geqslant 0}\in L^{1}(B_{t})\,|\,M_{t}=M_{0}+\int_{0}^{t}\sigma_{s}\,\mathrm{d}B_{s}\implies\right.$
$\hskip 145.10922ptM_{0}\sim\mu,M_{1}\sim\nu,(M_{t})\text{ is a martingale}\Bigg{\}},$ (1.1)
where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion and $(\mu,\nu)$ are convexly ordered probability measures with finite second moments. If $(\mu,\nu)$ is irreducible in the sense of [2], Appendix A.1, the solutions to the MBBP are called $(\mu,\nu)$-stretched Brownian motions, and have the form $\mathbb{E}[F(B_{1}+Y_{0})\,|\,\mathcal{F}_{t}]$, where $(\mathcal{F}_{t})$ is the filtration generated by $(B_{t}+Y_{0})$. The function $F:\mathbb{R}\rightarrow\mathbb{R}$ and the random variable $Y_{0}$ must be selected such that it is independent of $(B_{t})_{t>0}$, $F(B_{1}+Y_{0})\sim\nu$, and $\mathbb{E}[F(B_{1}+Y_{0})\,|\,Y_{0}]\sim\mu$. In the case where $(\mu,\nu)$ have irreducible components, the solution to the MBBP is a stretched Brownian motion on each irreducible component, see Theorem 3.1 in [1]. Stretched Brownian motions are interpolating martingales between the given measures $\mu$ and $\nu$. Another weak solution to the martingale interpolation problem is given by Schrödinger’s volatility models (SVMs) [10]. Inspired by the entropic relaxation of optimal transport ([18], [24]), also known as Schrödinger’s problem in specific cases [27], SVMs rely on a measure change. Consider two, possibly correlated, Brownian motions $(B_{t})_{t\geqslant 0}$ and $(\tilde{B}_{t})_{t\geqslant 0}$, under a measure $\mathbb{P}$, along with the system of SDEs on a real interval $[T_{0},T_{1}]$
$\,\mathrm{d}M_{t}=a_{t}\,M_{t}\,\mathrm{d}\widetilde{B}_{t},$
$\,\mathrm{d}a_{t}=b(a_{t})+c\left(a_{t}\right)\text{d}B_{t},$ (1.2)
where the real functions $b$ and $c$ are given, and are such that $(a_{t})_{t\in[T_{0},T_{1}]}$ is an Itô process. The goal is to change the measure $\mathbb{P}$ to an equivalent one, such that the process $(M_{t})_{0\leqslant t\leqslant 1}$ matches the given measures $\mu$ and $\nu$ at times zero and one, respectively, while being a true martingale. If several equivalent measures exist that satisfy these constraints, one must select the one that is closest to $\mathbb{P}$ in terms of the Kullback-Leibler divergence. By Girsanov’s theorem, changing $\mathbb{P}$ to an equivalent measure $\mathbb{Q}$ transforms the system of SDEs as follows:
$\,\mathrm{d}M_{t}=a_{t}M_{t}(\text{d}\widetilde{B}_{t}+\widetilde{\lambda}_{t}\,\mathrm{d}t),$
$\,\mathrm{d}a_{t}=b(a_{t})+c\left(a_{t}\right)(\text{d}B_{t}+\lambda_{t}\,\mathrm{d}t).$ (1.3)
One must then find the drifts $(\tilde{\lambda}_{t})$ and $(\lambda_{t})$, generated by the same measure change, that minimize the functional $\mathbb{E}_{\mathbb{Q}}[\int_{T_{0}}^{T_{1}}\widetilde{\lambda}^{2}_{s}+\lambda^{2}_{s}\,\mathrm{d}s]$ while satisfying the constraints $M_{0}\sim\mu$ and $M_{1}\sim\nu$, and so that $(M_{t})_{t\in[T_{0},T_{1}]}$ be a martingale.
In this paper, inspired by the information-based approach, see [7] and [8], we construct a class of martingales that match, almost surely, any set of convexly ordered random variables, indexed by an arbitrary set of fixed times. The construction of such interpolating martingales, called filtered arcade martingales (FAMs), need two main ingredients: a discrete-time martingale (for example, a solution to an MOT problem) and an interpolating process that we call the randomized arcade process (RAP).
In Section 2, we first develop the arcade processes (AP). Defined as a functional of a given stochastic process, called the driver, APs are sample-continuous stochastic processes that interpolate between zeros on the whole probability space, i.e., omega by omega. To achieve the interpolation, APs rely on deterministic functions, the interpolating coefficients. APs may be viewed as multi-period anticipative stochastic bridges. We study their properties and focus on Gaussian APs, which play an important role in the construction of FAMs. Starting from a Gauss-Markov driver, we show that it is always possible to construct a Markovian AP by utilizing the covariance structure of the driver. The resulting AP from this procedure is called a standard AP. This process is not unique, since one can construct infinitely many Markovian APs that are driven by the same Gauss-Markov process.
Section 3 treats an additive randomization of APs, that is, an AP added to a stochastic process that interpolates deterministically between the given random variables. We term such processes randomized arcade processes (RAPs), which can be thought of as a sum of a signal function and a noise process. By construction, a RAP can match any random variables on the whole probability space at an arbitrary sequence of fixed times, i.e, it is a stochastic interpolator between target random variables in the strong sense. The notion of the Markov property does not suit RAPs in general, since their natural filtration contains the $\sigma$-algebras generated by each previously matched target random variables. For this reason, we introduce a weaker notion, the conditionally-Markov property, and go on to show under which conditions a RAP is conditionally-Markov.
In the fourth section, we introduce the filtered arcade martingales (FAMs). A FAM is defined by the conditional expectation of the final target random variable, given the information generated by a RAP and, hence, inherits the filtering framework of the information-based approach that appears in [19], [7], [8]: the signal is the final target random variable and it can only be observed through a noisy version, the RAP. FAMs are tractable, thanks to the conditionally-Markov property enjoyed by the underlying RAPs, and can be simulated using Bayes formula. Applying Itô’s formula, we derive the stochastic differential equation satisfied by a FAM, which reveals the structure of the underlying innovations process that is adapted to the filtration generated by the respective RAP. The same ideas are applied to introduce the filtered arcade reverse-martingales (FARMs), constructed using reverse-filtrations of RAPs. Finally, as an application of the this theory, we briefly discuss a similar problem to martingale optimal transport, that incorporates noise in the optimization process, inspired by the entropic regularization of optimal transport and Schrödinger’s problem. This approach selects the martingale coupling that maximizes the expectation of the cumulative cost in time between a target random variable and its associated FAM, for a given RAP.
In what follows, let $n\in\mathbb{N}_{0}$ and consider the collection of fixed dates $\{T_{i}\in\mathbb{R}\,|\,i=0,1,\ldots,n\}$ such that $0\leqslant T_{0}<T_{1}<T_{2}<\ldots<T_{n}<\infty$, and introduce the ordered sets
$\{T_{0},T_{n}\}_{*}$ $=$ $\{T_{0},T_{1},\ldots,T_{n}\},$
$(T_{0},T_{n})_{*}$ $=$ $\bigcup\limits_{i=0}^{n-1}(T_{i},T_{i+1}).$
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\left(D_{t}\right)_{t\in[T_{0},T_{n}]}$ a sample-continuous stochastic process such that $\mathbb{P}[D_{t}=0]<1$ whenever $t\in(T_{0},T_{n})_{*}$, and $X$ an $\mathbb{R}^{n+1}$-valued random vector that is independent of $(D_{t})$.
2 Arcade processes
We construct stochastic processes on $[T_{0},T_{n}]$, as a functional of $(D_{t})$, which match exactly 00 (for all $\omega\in\Omega$) at the given times $\{T_{0},T_{n}\}_{*}$. The first step is to introduce deterministic functions called interpolating coefficients.
Definition 2.1.
The functions $f_{0},f_{1},\ldots,f_{n}$ are interpolating coefficients on $\{T_{0},T_{n}\}_{*}$ if $f_{0},$ $f_{1},\ldots,f_{n}\in C^{0}\left([T_{0},T_{n}],\mathbb{R}\right),f_{i}(T_{i})=1,$ and $f_{i}(T_{j})=0\text{ for }i,j=0,\ldots,n,\,i\neq j$.
We can now give the definition of what we call an arcade process.
Definition 2.2.
An arcade process (AP), denoted $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$, on the partition $\{T_{0},T_{n}\}_{*}$ is a stochastic process of the form
$A_{t}^{(n)}:=D_{t}-\sum_{i=0}^{n}f_{i}(t)D_{T_{i}},$ (2.1)
where $f_{0},\ldots,f_{n}$ are interpolating coefficients on $\{T_{0},T_{n}\}_{*}$. The process $(D_{t})_{t\in[T_{0},T_{n}]}$ is the driver of the AP. We denote by $(\mathcal{F}^{A}_{t})_{t\in[T_{0},T_{n}]}$ the filtration generated by $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$.
We observe that $A_{T_{0}}^{(n)}=A_{T_{1}}^{(n)}=\ldots=A_{T_{n}}^{(n)}=0$ by construction, for all $\omega\in\Omega$. That is, $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ strongly interpolates between zeros at times $t\in\{T_{0},T_{n}\}_{*}$.
Example 2.3.
For $n=1$, ${f_{0}(t)=\frac{t-T_{1}}{T_{0}-T_{1}}}$, ${f_{1}(t)=\frac{t-T_{0}}{T_{1}-T_{0}}}$, the AP driven by a standard Brownian motion $\left(B_{t}\right)_{t\geqslant 0}$ is the anticipative Brownian bridge on $[T_{0},T_{1}]$,
$A_{t}^{(1)}=B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}B_{T_{0}}-\frac{t-T_{0}}{T_{1}-T_{0}}B_{T_{1}}.$ (2.2)
Example 2.4.
For $n>1$, we can generalize the anticipative Brownian bridge by taking
$f_{0}(t)=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\frac{t-T_{n-1}}{T_{n}-T_{n-1}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (2.3)
and
$f_{i}(t)=\frac{t-T_{i-1}}{T_{i}-T_{i-1}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\frac{T_{i+1}-t}{T_{i+1}-T_{i}}\mathds{1}_{(T_{i},T_{i+1}]}(t),\quad\text{for }i=1,\ldots,n-1.$ (2.4)
We call this AP the stitched Brownian AP, for it can be written as
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}B_{T_{0}}-\frac{t-T_{0}}{T_{1}-T_{0}}B_{T_{1}},&\mbox{if }t\in[T_{0},T_{1}),\\ \\ B_{t}-\frac{T_{2}-t}{T_{2}-T_{1}}B_{T_{1}}-\frac{t-T_{1}}{T_{2}-T_{1}}B_{T_{2}},&\mbox{if }t\in[T_{1},T_{2}),\\ \,\vdots\\ B_{t}-\frac{T_{n}-t}{T_{n}-T_{n-1}}B_{T_{n-1}}-\frac{t-T_{n-1}}{T_{n}-T_{n-1}}B_{T_{n}},&\mbox{if }t\in[T_{n-1},T_{n}].\end{array}\right.$ (2.5)
Figure 1: Paths simulation of a stitched Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.4.
Example 2.5.
Another way of generalizing the anticipative Brownian bridge to obtain an AP driven by Brownian motion is by using Lagrange’s polynomial interpolation, that is
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}\quad\text{for }i=0,\ldots,n.$ (2.6)
We may call the resulting AP the Lagrange-Brownian AP, which has the form
$A_{t}^{(n)}=B_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}B_{T_{i}}.$ (2.7)
More generally, the Lagrange AP driven by a stochastic process $(D_{t})$ has the form
$A_{t}^{(n)}=D_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}D_{T_{i}}.$ (2.8)
Figure 2: Paths simulation of a Lagrange-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.5.
The Lagrange APs inherit Runge’s phenomenon from their interpolation coefficients. When $n$ is large, the AP oscillates around the edges of the interval.
Figure 3: Paths simulation of a Lagrange-Brownian AP with $n=10$, using the equidistant partition $\{T_{i}=i\,|\,i=0,1,\ldots,10\}$, illustrating Runge’s phenomenon, c.f., Example 2.5.
One can control this effect by applying a transformation to the interpolating coefficients. For instance, the map $x\mapsto|x|^{2(1-|x|)}$ applied to each of the interpolating coefficients
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}$
for $i=0,\ldots,n$, yields another set of interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{n}$ which do no longer suffer from Runge’s phenomenon.
Figure 4: Paths simulation of a Brownian AP with $n=10$ and interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{10}$, using the equidistant partition $\{T_{i}=i\,|\,i=0,1,\ldots,10\}$, c.f., Example 2.5.
Example 2.6.
Elliptic APs have interpolating coefficients given by
$f_{0}(t)=\sqrt{1-\left(\frac{t-T_{0}}{T_{1}-T_{0}}\right)^{2}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\sqrt{1-\left(\frac{t-T_{n}}{T_{n}-T_{n-1}}\right)^{2}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (2.9)
$f_{i}(t)=\sqrt{1-\left(\frac{t-T_{i}}{T_{i}-T_{i-1}}\right)^{2}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\sqrt{1-\left(\frac{t-T_{i}}{T_{i+1}-T_{i}}\right)^{2}}\mathds{1}_{(T_{i},T_{i+1}]}(t),\quad\text{for }i=1,\ldots,n-1.$ (2.10)
Figure 5: Paths simulation of an elliptic-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.6.
The expectation and the covariance of an AP, when they exist, are fully characterised by the driver and the interpolating coefficients.
Proposition 2.7.
If the driver $(D_{t})$ has a mean function $\mu_{D}$, a variance function $\sigma_{D}^{2}$, and a covariance function $K_{D}$, we have:
$\mu_{A}(t):=\mathbb{E}[A_{t}^{(n)}]=\mu_{D}(t)-\sum\limits_{i=0}^{n}f_{i}(t)\mu_{D}(T_{i}),$ (2.11)
$\sigma_{A}^{2}(t):=\mathrm{Var}{[A_{t}^{(n)}]}=\sigma_{D}^{2}(t)+\sum\limits_{i=0}^{n}f_{i}^{2}(t)\sigma_{D}^{2}(T_{i})-2f_{i}(t)K_{D}(t,T_{i})$
$\hskip 44.10185pt+2\sum\limits_{i=0}^{n}\sum\limits_{j=i+1}^{n}f_{i}(t)f_{j}(t)K_{D}(T_{i},T_{j}),$ (2.12)
$K_{A}(s,t):=\mathrm{Cov}[{A_{s}^{(n)},A_{t}^{(n)}}]=K_{D}(s,t)-\sum\limits_{i=0}^{n}f_{i}(t)K_{D}(s,T_{i})+f_{i}(s)K_{D}(t,T_{i})$
$\hskip 44.10185pt+\sum\limits_{i=0}^{n}\sum\limits_{j=0}^{n}f_{i}(s)f_{j}(t)K_{D}(T_{i},T_{j}).$ (2.13)
Proof.
Eq. (2.11) follows from the linearity property of the expectation and Eq. (2.12) follows from Eq. (2.13). Hence, it is enough to prove Eq. (2.13).
$\mathrm{Cov}[{A_{s}^{(n)},A_{t}^{(n)}}]$
$=\mathrm{Cov}\left[D_{s}-\sum_{i=0}^{n}f_{i}(s)D_{T_{i}},D_{t}-\sum_{i=0}^{n}f_{i}(t)D_{T_{i}}\right]$ (2.14)
$=K_{D}(s,t)-\mathrm{Cov}\left[D_{s},\sum_{i=0}^{n}f_{i}(t)D_{T_{i}}\right]-\mathrm{Cov}\left[\sum_{i=0}^{n}f_{i}(s)D_{T_{i}},D_{t}\right]$
$\hskip 190.91818pt+\mathrm{Cov}\left[\sum_{i=0}^{n}f_{i}(s)D_{T_{i}},\sum_{i=0}^{n}f_{i}(t)D_{T_{i}}\right]$ (2.15)
$=K_{D}(s,t)-\sum_{i=0}^{n}f_{i}(t)K_{D}(s,T_{i})-\sum_{i=0}^{n}f_{i}(s)K_{D}(T_{i},t)+\sum_{i=0}^{n}\sum_{j=0}^{n}f_{i}(s)f_{j}(t)K_{D}(T_{i},T_{j})$ (2.16)
$=K_{D}(s,t)-\sum\limits_{i=0}^{n}f_{i}(t)K_{D}(s,T_{i})+f_{i}(s)K_{D}(t,T_{i})+\sum\limits_{i=0}^{n}\sum\limits_{j=0}^{n}f_{i}(s)f_{j}(t)K_{D}(T_{i},T_{j}).$ (2.17)
Markov APs will play a crucial role in the construction of filtered arcade martingales. A useful property of these processes is the following:
Proposition 2.8.
Let $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ be an AP on $\{T_{0},T_{n}\}_{*}$ that is Markov with respect to $(\mathcal{F}^{A}_{t})_{t\in[T_{0},T_{n}]}$. If $t>T_{i}$ for $T_{i}\in\{T_{0},T_{n}\}_{*}$, then $A_{t}^{(n)}$ is independent of $\mathcal{F}^{A}_{T_{i}}$.
In particular, if $(A_{t}^{(n)})$ is Markov and has a covariance function, then $\mathrm{Cov}(A_{s}^{(n)},A_{t}^{(n)})=0$ whenever $s\in[T_{i},T_{i+1}]$, $t\in[T_{j},T_{j+1}]$ and $j\neq i$.
If we think of APs as noise processes, it is natural to study their Gaussian subclass. But instead of looking at all APs that are also Gaussian processes, we focus on APs driven by Gaussian processes.
Definition 2.9.
An AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is said to be a Gaussian AP if its driver $(D_{t})$ is a Gaussian stochastic process on $[T_{0},T_{n}]$.
Not all APs that are Gaussian processes are Gaussian APs, though all Gaussian APs are Gaussian processes. For example, take $D_{t}=B_{t}+tY$, where $(B_{t})$ is a Gaussian process and $Y$ a non-Gaussian random variable, and $A_{t}=D_{t}-(1-t)D_{0}-tD_{1}$. The driver is not Gaussian while the AP is a Gaussian process. Next, we give a first result on the Markov property of Gaussian APs.
Theorem 2.10.
Let $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ be a Gaussian AP on $\{T_{0},T_{n}\}_{*}$ with covariance function $K_{A}$. Then $(A_{t}^{(n)})$ is Markov with respect to its own filtration if and only if for all $(r,s,t)\in(T_{0},T_{n})_{*}^{3}$ satisfying $r\leqslant s<t$, there exists $a(s,t)\in\mathbb{R}_{0}$ such that
$K_{A}(r,t)=\left\{\begin{array}[]{cc}0,&\mbox{ if }t\in[T_{i},T_{i+1}],r\in[T_{j},T_{j+1}],i\neq j,\\ \\ K_{A}(r,s)a(s,t),&\mbox{ otherwise}.\end{array}\right.$ (2.18)
Proof.
Suppose $K_{A}$ is of the form 2.18, we shall show that $(A_{t}^{(n)})$ is Markov. Let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(A_{t}^{(n)})$ is Markov if and only if
$\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{1}}^{(n)},\ldots,A_{s_{k}}^{(n)}\right]=\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{k}}^{(n)}\right].$ (2.19)
Since $K_{A}(s,t)=0$ unless $s$ and $t$ are in the same sub-interval, we can assume without loss of generality that $(s_{1},s_{2},\ldots,s_{k},t)\in(T_{m},T_{m+1})^{k+1}$ for a fixed $m\in\{0,\ldots,n-1\}$. Define
$\Delta_{q}=\sum_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)},\quad q=1,\ldots,k-1,$ (2.20)
where the coefficients $(c_{i,q})$ are real numbers satisfying
$\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0\quad\mbox{ and }\quad\mathbb{\det}\begin{pmatrix}c_{1,1}&\ldots&c_{1,k-1}\\ \vdots&&\vdots\\ c_{k-1,1}&\ldots&c_{k-1,k-1}\\ \end{pmatrix}\neq 0.$ (2.21)
We notice that
$\sigma(A_{s_{1}}^{(n)},\ldots,A_{s_{k}}^{(n)})=\sigma(\Delta_{1},\ldots,\Delta_{k-1},A_{s_{k}}^{(n)})\iff\mathbb{\det}\begin{pmatrix}c_{1,1}&\ldots&c_{1,k-1}\\ \vdots&&\vdots\\ c_{k-1,1}&\ldots&c_{k-1,k-1}\\ \end{pmatrix}\neq 0.$ (2.22)
It remains to be shown that $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1}).$ Equivalently, because we are treating the Gaussian case, we shall show that $\mathrm{Cov}[A_{t}^{(n)},\Delta_{q}]=0\mbox{ and }\mathrm{Cov}[A_{s_{k}}^{(n)},\Delta_{q}]=0\mbox{ for }q=1,\ldots,k-1.$ Expanding these covariances, we get
$\mathrm{Cov}[A_{t}^{(n)},\Delta_{q}]=\sum_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0,$ (2.23)
which is guaranteed by Eq. (2.21), and
$\mathrm{Cov}[A_{s_{k}}^{(n)},\Delta_{q}]=\sum_{i=1}^{k}c_{i,q}K_{A}(s_{k},s_{i})=0.$ (2.24)
The first equation implies the second because $K_{A}(s_{i},t)=0\implies K_{A}(s_{k},s_{i})=0$ by 2.18, and, when $K_{A}(s_{i},t)\neq 0$, we have $K_{A}(s_{i},t)=a(s_{k},t)K_{A}(s_{k},s_{i})$, so that
$\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0\implies\sum\limits_{i=1}^{k}c_{i,q}a(s_{k},t)K_{A}(s_{k},s_{i})=0\implies\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{k},s_{i})=0.$ (2.25)
This concludes one implication.
For the converse, suppose without loss of generality that the driver has mean 00. We observe that $K_{A}(x,y)=0$ if $x\in[T_{i},T_{i+1}],y\in[T_{j},T_{j+1}],i\neq j$. Let $(r,s,t)\in(T_{i},T_{i+1})^{3}$ such that $r\leqslant s<t$. Since $(A_{t}^{(n)})$ is Gaussian, we have
$\mathbb{E}[A_{t}^{(n)}\,|\,A_{s}^{(n)}]=\frac{K_{A}(s,t)}{K_{A}(s,s)}A_{s}^{(n)}.$ (2.26)
Using the Markov property of $(A_{t}^{(n)})$, we get
$K_{A}(r,t)=\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}\,|\,A_{s}^{(n)}]]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}\,|\,A_{s}^{(n)}]\mathbb{E}[A_{t}^{(n)}\,|\,A_{s}^{(n)}]],$ (2.27)
which implies
$K_{A}(r,t)=\mathbb{E}\left[\frac{K_{A}(s,t)}{K_{A}(s,s)}A_{s}^{(n)}\frac{K_{A}(r,s)}{K_{A}(s,s)}A_{s}^{(n)}\right]=\frac{K_{A}(s,t)K_{A}(r,s)}{K_{A}(s,s)}.$ (2.28)
Hence ${a(s,t)=\frac{K_{A}(s,t)}{K_{A}(s,s)}}$. ∎
We can simplify the statement of Theorem 2.10 in the following way.
Corollary 2.11.
A Gaussian AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is Markov with respect to its own filtration if and only if there exist real functions $A_{1}:[T_{0},T_{n}]\rightarrow\mathbb{R}$ and $A_{2}:[T_{0},T_{n}]\rightarrow\mathbb{R}$ such that
$K_{A}(s,t)=\sum_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t).$ (2.29)
Proof.
If
$K_{A}(s,t)=\sum_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t),$ (2.30)
then $\forall\,(r,s,t)\in(T_{0},T_{n})_{*}^{3}$ such that $r\leqslant s<t$, one has
$K_{A}(r,t)=\left\{\begin{array}[]{cc}0,&\mbox{ if }t\in[T_{i},T_{i+1}],r\in[T_{j},T_{j+1}],i\neq j,\\ \\ K_{A}(r,s)\frac{A_{2}(t)}{A_{2}(s)},&\mbox{ otherwise}.\end{array}\right.$ (2.31)
Hence, $(A_{t}^{(n)})$ is Markov.
Conversly, suppose that $(A_{t}^{(n)})$ is Markov, and let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$. Then,
$\frac{K_{A}(r,t)}{K_{A}\left(r,\frac{T_{m}+T_{m+1}}{2}\right)}=\frac{K_{A}\left(\frac{T_{m}+T_{m+1}}{2},t\right)}{K_{A}\left(\frac{T_{m}+T_{m+1}}{2},\frac{T_{m}+T_{m+1}}{2}\right)}$ (2.32)
for any $(r,t)\in(T_{m},T_{m+1})^{2}$ such that $r<\frac{T_{m}+T_{m+1}}{2}<t$ by Theorem 2.10. Hence, if
$A_{1}(x)=\sum_{i=0}^{n-1}K_{A}\left(x,\frac{T_{i}+T_{i+1}}{2}\right)\mathds{1}_{(T_{i},T_{i+1})}(x),$ (2.33)
and
$A_{2}(x)=\sum_{i=0}^{n-1}\frac{K_{A}\left(\frac{T_{i}+T_{i+1}}{2},x\right)}{K_{A}\left(\frac{T_{i}+T_{i+1}}{2},\frac{T_{i}+T_{i+1}}{2}\right)}\mathds{1}_{(T_{i},T_{i+1})}(x),$ (2.34)
we have
$K_{A}(s,t)=\sum_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t).$ (2.35)
If $T_{i}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{i},T_{i+1})^{2}$ such that $s<t$, then $\lim\limits_{s\rightarrow T_{i}}A_{1}(s)=0$ or $\lim\limits_{t\rightarrow T_{i+1}}A_{2}(t)=0$ by continuity of $K_{A}$, and $A_{1}/A_{2}(t)$ is positive and non-decreasing on each interval $(T_{i},T_{i+1})$, since $K_{A}$ is a covariance function.
Starting from a Gauss-Markov driver $(D_{t})$, it is always possible to construct a Markovian $(A_{t}^{(n)})$ by applying the following procedure.
Theorem 2.12.
For any Gauss-Markov driver $(D_{t})_{t\in[T_{0},T_{n}]}$, there exists an AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$, driven by $(D_{t})$, that is Markovian.
Proof.
Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ such that $s<t$. Choose the interpolating coefficients $f_{0},\ldots,f_{n}$ such that
$\begin{pmatrix}K_{D}(T_{0},T_{0})&\ldots&K_{D}(T_{0},T_{n})\\ \vdots&\ddots&\vdots\\ K_{D}(T_{n},T_{0})&\ldots&K_{D}(T_{n},T_{n})\end{pmatrix}\begin{pmatrix}f_{0}(\cdot)\\ \vdots\\ f_{n}(\cdot)\end{pmatrix}=\begin{pmatrix}K_{D}(\cdot,T_{0})\\ \vdots\\ K_{D}(\cdot,T_{n})\end{pmatrix}.$ (2.36)
Then,
$\sum_{j=0}^{n}f_{j}(\cdot)K_{D}(T_{i},T_{j})-K_{D}(\cdot,T_{i})=0,\quad\forall i=0,\ldots,n,$ (2.37)
which implies
$K_{A}(x,y)=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(x)K_{D}(y,T_{i})=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(y)K_{D}(x,T_{i}),\quad\forall(x,y)\in[T_{0},T_{n}]^{2}.$ (2.38)
Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ such that $s<t$. Recalling that $(D_{t})$ is Gauss-Markov, there exist two functions, $H_{1}:(T_{m},T_{m+1})\rightarrow\mathbb{R}$ and $H_{2}:(T_{m},T_{m+1})\rightarrow\mathbb{R}$, such that $K_{D}(s,t)=H_{1}(s)H_{2}(t)$. So, one can write
$K_{A}(s,t)$ $=H_{1}(s)H_{2}(t)-\sum_{i=0}^{m}f_{i}(s)H_{1}(T_{i})H_{2}(t)-\sum_{i=m+1}^{n}f_{i}(s)H_{1}(t)H_{2}(T_{i})$
$=H_{2}(t)\left(H_{1}(s)-\sum_{i=0}^{m}f_{i}(s)H_{1}(T_{i})\right)-H_{1}(t)\sum_{i=m+1}^{n}f_{i}(s)H_{2}(T_{i})$
$=\left(\sum_{i=m+1}^{n}f_{i}(s)H_{2}(T_{i})\right)\left(\frac{H_{1}(T_{m+1})}{H_{2}(T_{m+1})}H_{2}(t)-H_{1}(t)\right)$ (2.39)
where we used Eq. (2.37) with $i=m+1$. Hence,
$A_{1}(x)\mathds{1}_{(T_{m},T_{m+1})}(x)$ $=\left(\sum_{i=m+1}^{n}f_{i}(x)H_{2}(T_{i})\right)\mathds{1}_{(T_{m},T_{m+1})}(x),$ (2.40)
$A_{2}(x)\mathds{1}_{(T_{m},T_{m+1})}(x)$ $=\left(\frac{H_{1}(T_{m+1})}{H_{2}(T_{m+1})}H_{2}(x)-H_{1}(x)\right)\mathds{1}_{(T_{m},T_{m+1})}(x),$ (2.41)
and $(A_{t}^{(n)})$ is Markov by Corollary 2.11. ∎
Recall that we imposed on all drivers $(D_{t})$ the property that $\mathbb{P}[D_{t}=0]<1$ for $t\in(T_{0},T_{n})_{*}$. If we extend this property to $[T_{0},T_{n}]$ instead, then the above construction of a Markovian AP becomes explicit.
Corollary 2.13.
If $(D_{t})$ is a Gauss-Markov process such that $\mathbb{P}[D_{t}=0]<1$ for $t\in[T_{0},T_{n}]$, where $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$ for some real functions $H_{1}$ and $H_{2}$, then the solution to Eq. (2.37) is given by
$f_{0}(x)$ $=\frac{H_{1}(T_{1})H_{2}(x)-H_{1}(x)H_{2}(T_{1})}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}\mathds{1}_{[T_{0},T_{1}]}(x),$ (2.42)
$f_{i}(x)$ $=\frac{H_{1}(x)H_{2}(T_{i-1})-H_{1}(T_{i-1})H_{2}(x)}{H_{1}(T_{i})H_{2}(T_{i-1})-H_{1}(T_{i-1})H_{2}(T_{i})}\mathds{1}_{[T_{i-1},T_{i}]}(x)$
$\quad+\frac{H_{1}(T_{i+1})H_{2}(x)-H_{1}(x)H_{2}(T_{i+1})}{H_{1}(T_{i+1})H_{2}(T_{i})-H_{1}(T_{i})H_{2}(T_{i+1})}\mathds{1}_{(T_{i},T_{i+1}]}(x),\quad\text{for }i=1,\ldots,n-1,$ (2.43)
$f_{n}(x)$ $=\frac{H_{1}(x)H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(x)}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}\mathds{1}_{(T_{n-1},T_{n}]}(x).$ (2.44)
Proof.
Let ($T_{m^{-}},T_{m},T_{m^{+}})\in\{T_{0},T_{n-1}\}_{*}^{3}$ such that $T_{m^{-}}\leqslant T_{m}<T_{m^{+}}$, and $x\in(T_{m},T_{m+1})$. Then,
$\sum_{j=0}^{n}f_{j}(x)K_{D}(T_{m^{-}},T_{j})$ $=f_{m}(x)H_{1}(T_{m^{-}})H_{2}(T_{m})+f_{m+1}(x)H_{1}(T_{m^{-}})H_{2}(T_{m})$
$=\frac{H_{1}(T_{m+1})H_{2}(x)-H_{1}(x)H_{2}(T_{m+1})}{H_{1}(T_{m+1})H_{2}(T_{m})-H_{1}(T_{m})H_{2}(T_{m+1})}H_{1}(T_{m^{-}})H_{2}(T_{m})$
$\hskip 31.2982pt+\frac{H_{1}(x)H_{2}(T_{m})-H_{1}(T_{m})H_{2}(x)}{H_{1}(T_{m+1})H_{2}(T_{m})-H_{1}(T_{m})H_{2}(T_{m+1})}H_{1}(T_{m^{-}})H_{2}(T_{m+1})$
$=\frac{H_{1}(x)(H_{1}(T_{m^{-}})H_{2}(T_{m})H_{2}(T_{m+1})-H_{1}(T_{m^{-}})H_{2}(T_{m})H_{2}(T_{m+1}))}{H_{1}(T_{m+1})H_{2}(T_{m})-H_{1}(T_{m})H_{2}(T_{m+1})}$
$\hskip 31.2982pt+\frac{H_{2}(x)(H_{1}(T_{m+1})H_{1}(T_{m^{-}})H_{2}(T_{m})-H_{1}(T_{m})H_{1}(T_{m^{-}})H_{2}(T_{m+1}))}{H_{1}(T_{m+1})H_{2}(T_{m})-H_{1}(T_{m})H_{2}(T_{m+1})}$
$=H_{1}(T_{m^{-}})H_{2}(x)=K_{D}(T_{m^{-}},x)$ (2.45)
The same argument applies to show $\sum_{j=0}^{n}f_{j}(x)K_{D}(T_{m^{+}},T_{j})=K_{D}(x,T_{m^{+}})$. Hence,
$\begin{pmatrix}K_{D}(T_{0},T_{0})&\ldots&K_{D}(T_{0},T_{n})\\ \vdots&\ddots&\vdots\\ K_{D}(T_{n},T_{0})&\ldots&K_{D}(T_{n},T_{n})\end{pmatrix}\begin{pmatrix}f_{0}(\cdot)\\ \vdots\\ f_{n}(\cdot)\end{pmatrix}=\begin{pmatrix}K_{D}(\cdot,T_{0})\\ \vdots\\ K_{D}(\cdot,T_{n})\end{pmatrix}.$ (2.46)
Next, recalling the steps following Eq. (2.37) concludes the proof. ∎
Remark 2.14.
Notice that if $(D_{t})$ does not satisfy $\mathbb{P}[D_{t}=0]<1$ for some $t\in\{T_{0},T_{n}\}_{*}$, it is still straightforward to construct the above AP by removing all the rows and columns of zeros in the matrix
$\begin{pmatrix}K_{D}(T_{0},T_{0})&\ldots&K_{D}(T_{0},T_{n})\\ \vdots&\ddots&\vdots\\ K_{D}(T_{n},T_{0})&\ldots&K_{D}(T_{n},T_{n})\end{pmatrix}.$ (2.47)
Since the driver already matches 00 at $t\in\{T_{0},T_{n}\}_{*}$, there is no need to find the corresponding interpolating coefficients because they do not appear in the AP expression. This is illustrated by taking the Brownian motion as a driver with $T_{0}=0$. Then $f_{0}$ does not matter since it is multiplied by 00 in the AP expression.
Remark 2.15.
Inserting the interpolating coefficients given in Corollary 2.13 into Eqs. (2.40) and (2.41) yields
$A_{1}(x)\mathds{1}_{(T_{m},T_{m+1})}(x)$ $=\frac{H_{1}(x)H_{2}(T_{m})-H_{1}(T_{m})H_{2}(x)}{H_{1}(T_{m+1})H_{2}(T_{m})-H_{1}(T_{m})H_{2}(T_{m+1})}H_{2}(T_{m+1})\mathds{1}_{(T_{m},T_{m+1})}(x),$ (2.48)
$A_{2}(x)\mathds{1}_{(T_{m},T_{m+1})}(x)$ $=\left(\frac{H_{1}(T_{m+1})}{H_{2}(T_{m+1})}H_{2}(x)-H_{1}(x)\right)\mathds{1}_{(T_{m},T_{m+1})}(x),$ (2.49)
for $m=1,\ldots,n-1$.
This method to construct Markovian APs is not unique, but can be perceived as natural. The resulting APs are called standard.
Definition 2.16.
A standard AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is an AP, driven by a Gauss-Markov process $(D_{t})$, of the form
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}D_{t}-\frac{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{0}}-\frac{H_{1}(t)H_{2}(T_{0})-H_{1}(T_{0})H_{2}(t)}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{1}}&\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{H_{1}(T_{2})H_{2}(t)-H_{1}(t)H_{2}(T_{2})}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{1}}-\frac{H_{1}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}(t)}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{2}}&\mbox{for }t\in[T_{1},T_{2}),\\ \,\vdots\\ D_{t}-\frac{H_{1}(T_{n})H_{2}(t)-H_{1}(t)H_{2}(T_{n})}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n-1}}&\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{H_{1}(t)H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(t)}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n}}&\mbox{for }t\in[T_{n-1},T_{n}],\end{array}\right.$ (2.50)
where $H_{1}$ and $H_{2}$ are the real functions appearing in the covariance function of the driver: $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$.
Example 2.17.
If $(D_{t})$ is an Ornstein-Uhlenbeck process with parameters $\theta>0,\sigma>0,\mu\in\mathbb{R}$ and initial value $d_{0}\in\mathbb{R}$, that is, the solution to
$\,\mathrm{d}D_{t}=\theta\left(\mu-D_{t}\right)\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},\quad D_{0}=d_{0},$ (2.51)
then $K_{D}(s,t)=\sigma^{2}/(2\theta)\exp{[\theta\min(s,t)]}\exp{[-\theta\max(s,t)]}$. The standard AP driven by $(D_{t})$ is
$A_{t}^{(n)}=\left\{\begin{array}[]{l}D_{t}-\frac{\exp{[\theta(T_{1}-t)]}-\exp{[-\theta(T_{1}-t)]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{0}}-\frac{\exp{[\theta(t-T_{0})]}-\exp{[-\theta(t-T_{0})]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{1}}\quad\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{\exp{[\theta(T_{2}-t)]}-\exp{[-\theta(T_{2}-t)]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{1}}-\frac{\exp{[\theta(t-T_{1})]}-\exp{[-\theta(t-T_{1})]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{2}}\quad\mbox{for }t\in[T_{1},T_{2}),\\ \,\vdots\\ D_{t}-\frac{\exp{[\theta(T_{n}-t)]}-\exp{[-\theta(T_{n}-t)]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n-1}}-\frac{\exp{[\theta(t-T_{n-1})]}-\exp{[-\theta(t-T_{n-1})]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n}}\\ \\ \hskip 324.36124pt\quad\mbox{for }t\in[T_{n-1},T_{n}].\end{array}\right.$ (2.52)
There are infinitely many Markovian APs driven by the same Gauss-Markov driver. In general, when $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ with $s<t$, we have
$K_{A}(s,t)=$ $\left(H_{1}(s)-\sum_{i=0}^{m}f_{i}(s)H_{1}(T_{i})\right)\left(H_{2}(t)-\sum_{i=m+1}^{n}f_{i}(t)H_{2}(T_{i})\right)$
$-\left(\sum_{i=m+1}^{n}f_{i}(s)H_{2}(T_{i})\right)\left(H_{1}(t)-\sum_{i=0}^{m+1}f_{i}(t)H_{1}(T_{i})\right)$
$+\left(\sum_{i=m+1}^{n}f_{i}(s)H_{1}(T_{i})\right)\sum_{i=m+2}^{n}f_{i}(t)H_{2}(T_{i})$
$-\left(H_{2}(s)-\sum_{i=0}^{m}f_{i}(s)H_{2}(T_{i})\right)\sum_{i=0}^{m}f_{i}(t)H_{1}(T_{i}),$ (2.53)
where we use the convention that an empty sum is equal to zero. There are as many Markovian APs driven by $(D_{t})$ as there are ways to express the covariance function $K_{A}(s,t)$ as a product of a function of $s$ and a function of $t$.
Example 2.18.
If $(D_{t})$ is a standard Brownian motion, applying Eq. (2.37) to find appropriate interpolating coefficients yields the stitched Brownian AP, see Example 2.4. But there are other Markovian APs driven by standard Brownian motion. For instance, in the two-period case, we may choose
$f_{0}(t)$ $=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{\left[T_{0},\frac{T_{1}+T_{2}}{2}\right]}(t)-\frac{T_{2}-t}{T_{1}-T_{0}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (2.54)
$f_{1}(t)$ $=\frac{t-T_{0}}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t)+\frac{T_{2}-t}{T_{2}-T_{1}}\mathds{1}_{(T_{1},T_{2}]}(t),$ (2.55)
$f_{2}(t)$ $=\frac{t-T_{1}}{T_{2}-T_{1}}\mathds{1}_{\left[T_{1},T_{2}\right]}(t).$ (2.56)
It is straightforward to verify that these are interpolating coefficients. Let $(A_{t}^{(2)})_{t\in[T_{0},T_{2}]}$ be the AP constructed with these interpolating coefficients and driven by a standard Brownian motion. Then,
$K_{A}(s,t)$
$=\frac{(\min(s,t)-T_{0})(T_{1}-\max(s,t))}{T_{1}-T_{0}}\mathds{1}_{(T_{0},T_{1})}(s,t)$
$+(\min(s,t)-T_{1})\left(\frac{T_{0}(\max(s,t)+T_{0})-3T_{0}T_{1}+T_{1}^{2}}{(T_{1}-T_{0})^{2}}+\frac{T_{1}-\max(s,t)}{T_{2}-T_{1}}\right)\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(s,t)$
$+\frac{(\min(s,t)-T_{1})(T_{2}-\max(s,t))}{T_{2}-T_{1}}\frac{T_{0}^{2}+T_{1}^{2}+T_{0}(T_{2}-3T_{1})}{(T_{1}-T_{0})^{2}}$
×$\hskip 241.84842pt\times\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(\min(s,t))\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(\max(s,t))$
$+\frac{\min(s,t)(T_{0}^{2}+T_{1}^{2}-T_{0}(T_{1}+T_{2}))-(T_{1}-T_{0})^{2}T_{1}+T_{0}T_{2}(T_{2}-T_{1})}{(T_{1}-T_{0})^{2}}\frac{T_{2}-\max(s,t)}{T_{2}-T_{1}}$
×$\hskip 241.84842pt\times\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(s,t).$ (2.57)
Thus,
$A_{1}(x)$ $=(x-T_{0})\mathds{1}_{\left[T_{0},T_{1}\right]}(x)+(x-T_{1})\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(x)$ (2.58)
$+\frac{x(T_{0}^{2}+T_{1}^{2}-T_{0}(T_{1}+T_{2}))-(T_{1}-T_{0})^{2}T_{1}+T_{0}T_{2}(T_{2}-T_{1})}{T_{0}^{2}+T_{1}^{2}+T_{0}(T_{2}-3T_{1})}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(x),$
$A_{2}(x)=\frac{T_{1}-x}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(x)$ $+\left(\frac{T_{0}(x+T_{0})-3T_{0}T_{1}+T_{1}^{2}}{(T_{1}-T_{0})^{2}}+\frac{T_{1}-x}{T_{2}-T_{1}}\right)\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(x)$
$+\frac{(T_{2}-x)}{T_{2}-T_{1}}\frac{T_{0}^{2}+T_{1}^{2}+T_{0}(T_{2}-3T_{1})}{(T_{1}-T_{0})^{2}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(x).$ (2.59)
Recalling Theorem 3.8, this AP is Markovian. This process is a slight modification of the stitched Brownian AP, where $f_{0}$ is not 00 on $(T_{1},T_{2})$. Hence, $B_{T_{0}}$ still has a negative influence on the paths of the AP on $(T_{1},T_{2})$, since $f_{0}$ is negative on $(T_{1},T_{2})$. Similarly, we can modify the stitched Brownian arcade by setting $f_{2}\neq 0$ on $(T_{0},T_{1})$:
$f_{0}(t)$ $=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t),$ (2.60)
$f_{1}(t)$ $=\frac{t-T_{0}}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t)+\frac{T_{2}-t}{T_{2}-T_{1}}\mathds{1}_{(T_{1},T_{2}]}(t),$ (2.61)
$f_{2}(t)$ $=\frac{T_{0}-t}{T_{2}-T_{1}}\mathds{1}_{\left[T_{0},\frac{T_{0}+T_{1}}{2}\right]}(t)+\frac{t-T_{1}}{T_{2}-T_{1}}\mathds{1}_{\left(\frac{T_{0}+T_{1}}{2},T_{2}\right]}(t).$ (2.62)
For this choice of interpolating coefficients, $B_{T_{2}}$ has an influence on the paths of the AP on $(T_{0},T_{1})$. Combining the interpolating coefficients $f_{0}$ from Eq. (2.54) and $f_{2}$ from Eq. (2.62), we can find an interpolating coefficient $f_{1}$, such that a Brownian AP with these interpolating coefficients is Markovian:
$f_{0}(t)$ $=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{\left[T_{0},\frac{T_{1}+T_{2}}{2}\right]}(t)+\frac{t-T_{2}}{T_{1}-T_{0}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (2.63)
$f_{1}(t)$ $=\frac{(t-T_{0})(T_{2}-T_{0})}{(T_{2}-T_{1})(T_{1}-T_{0})}\mathds{1}_{\left[T_{0},\frac{T_{0}+T_{1}}{2}\right]}(t)+\frac{T_{1}^{2}-T_{0}T_{2}+t(T_{0}-2T_{1}+T_{2})}{(T_{1}-T_{0})(T_{2}-T_{1})}\mathds{1}_{\left(\frac{T_{0}+T_{1}}{2},T_{1}\right]}(t)$
$+\frac{T_{1}(T_{0}(T_{1}-2T_{2})+T_{1}T_{2})+t(T_{0}T_{2}-T_{1}^{2})}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(t)$
$+\frac{(T_{2}-t)(T_{1}^{2}+T_{0}(T_{2}-2T_{1}))}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (2.64)
$f_{2}(t)$ $=\frac{T_{0}-t}{T_{2}-T_{1}}\mathds{1}_{\left[T_{0},\frac{T_{0}+T_{1}}{2}\right]}(t)+\frac{t-T_{1}}{T_{2}-T_{1}}\mathds{1}_{\left(\frac{T_{0}+T_{1}}{2},T_{2}\right]}(t).$ (2.65)
The key to building non-standard APs is to break each interval into several sub-intervals, and to define the interpolating coefficients piecewise on these sub-intervals while guaranteeing that they remain continuous, and that the expression of $K_{A}$ in Eq. (2.53) has separable variables.
3 Randomized arcade processes
We extend the construction of arcade processes to allow interpolation between the components of a random vector $X$ instead of interpolation between zeros. Two sets $\{f_{0},\ldots,f_{n}\}$ and $\{g_{0},\ldots,g_{n}\}$ of interpolating coefficients, see Definition 2.1, are needed to ensure the matching of the target random variables. We recall that the $\mathbb{R}^{n+1}$-valued random vector $X=(X_{0},\ldots,X_{n})$ is independent of the stochastic driver $(D_{t})$, while the random variables $X_{0},\ldots,X_{n}$ may be mutually dependent.
Definition 3.1.
An $X$-randomized arcade process (X-RAP), denoted $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ on the partition $\{T_{0},T_{n}\}_{*}$ is a stochastic process of the form
$I_{t}^{(n)}:=D_{t}-\sum_{i=0}^{n}\left(f_{i}(t)D_{T_{i}}-g_{i}(t)X_{i}\right),$ (3.1)
where $f_{0},\ldots,f_{n}$ and $g_{0},\ldots,g_{n}$ are interpolating coefficients on $\{T_{0},T_{n}\}_{*}$. Writing $I_{t}^{(n)}=S_{t}^{(n)}+A_{t}^{(n)}$, we refer to
$S_{t}^{(n)}=\sum\limits_{i=0}^{n}g_{i}(t)X_{i}$ (3.2)
as the signal function of $I_{t}^{(n)}$ and to
$A_{t}^{(n)}=D_{t}-\sum\limits_{i=0}^{n}f_{i}(t)D_{T_{i}}$ (3.3)
as the noise process of $I_{t}^{(n)}$. We denote by $(\mathcal{F}^{I}_{t})_{t\in[T_{0},T_{n}]}$ the filtration generated by $(I_{t}^{(n)})$.
We emphasis that $I_{T_{0}}^{(n)}=X_{0},\ldots,I_{T_{n}}^{(n)}=X_{n}$, so $(I_{t}^{(n)})$ is a strong stochastic interpolator between the random variables $X_{0},\ldots,X_{n}$. The signal function $(S_{t}^{(n)})$ is independent of the noise process $(A_{t}^{(n)})$ since $X$ is assumed independent of $(D_{t})_{t\in[T_{0},T_{n}]}$.
Remark 3.2.
A related but distinct class of processes, introduced in [22] and called random $n$-bridges, shares some characteristics with RAPs. These processes, defined weakly, match given probability measures instead of random variables. They are constructed in a similar fashion as stochastic bridges, that is, by conditioning a stochastic process to match given distributions at given times. In special cases, the law of a RAP satisfies the conditions for the RAP to be a random $n$-bridge. For instance, the RAP obtained by randomizing the stitched Brownian bridge, using the same interpolating coefficients for the signal function as the ones used in the noise process, has a law that satisfies the conditions for the RAP to be a random $n$-bridge. However, any other RAP driven by Brownian motion is not a random $n$-bridge. Conversely, certain random $n$-bridges cannot have the same law as a RAP. In this paper, we chose not to consider the case in which the driver underlying a RAP jumps. But if we did, just for the sake of comparison, a random $n$-bridge built using a gamma process will not match the law of any RAP driven by a gamma process, since APs are sums, not products. The development of alternative arcade processes driven by (pure) jump processes will be pursued in future research. It requires a different construction that does not rely on the covariance function, similar to the change needed to go from Gaussian bridges to Lévy bridges [12] or Markov [20] bridges.
The paths of an $X$-RAP depend on the joint distribution of the random vector $X$, not only on its marginal distributions. This fact is illustrated in the following example.
Example 3.3.
Let $X=(X_{0},\ldots,X_{5})$ be a vector of independent and uniform $\mathcal{U}\,(\{-1,1\})$ random variables, and $Y=(Y_{0},\ldots,Y_{5})$ be another vector of random variables such that $Y_{0}\sim\mathcal{U}\,(\{-1,1\})$, $Y_{i}=-Y_{i-1}$ for $i=1,\ldots,5$. Let $(A_{t}^{(5)})$ be an AP with elliptic interpolation coefficients driven by Brownian motion, $g_{i}=f_{i}$ for $i=0,\ldots,5$, and $(I_{t}^{(5)})$, $(\tilde{I}_{t}^{(5)})$ its associated $X$-RAP and $Y$-RAP, respectively. Although the same driver and interpolating coefficients for both RAPs are the same, and the vectors $X$ and $Y$ have the same marginal distributions, the paths of $(I_{t}^{(5)})$ and $(\tilde{I}_{t}^{(5)})$ are different.
Figure 6: Paths simulation of $(I_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 3.3.
Figure 7: Paths simulation of $(\tilde{I}_{t}^{(5)})$ on $[0,10]$ using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 3.3.
Besides its main purpose of interpolating in the strong sense, a RAP can also be used to mimic another stochastic process. Let $(Y_{t})_{t\in[T_{0},T_{n}]}$ be a sample-continuous stochastic process. For instance, if $\{T_{0},T_{n}\}_{*}$ is the equidistant partition of an interval $[a,b]$, $X=(Y_{T_{0}},\ldots,Y_{T_{n}})$, and $\{f_{0},\ldots,f_{n}\}=\{g_{0},\ldots,g_{n}\}$ are the piecewise linear interpolating coefficients, then for almost all $\omega\in\Omega$, $\sup\limits_{t\in[a,b]}A_{t}^{(n)}(\omega)\rightarrow 0$, and $\sup\limits_{t\in[a,b]}S_{t}^{(n)}(\omega)\rightarrow Y_{t}(\omega)$ as $n\rightarrow\infty$. Hence, in that case,
[…]
with probability one.
Example 3.4.
Let $[a,b]=[0,10]$, and $\{T_{0},T_{n}\}_{*}$ be its equidistant partition. If $(Y_{t})$ is a fractional Brownian motion (fBm) and $X=(Y_{T_{0}},\ldots,Y_{T_{n}})$, then the paths of an $X$-randomized stitched Brownian arcade are similar to the one of $(Y_{t})$ when $n$ is large enough.
Figure 8: Path of the fBm $(Y_{t})$ , c.f., Example 3.4.
Figure 9: Path of an $X$-RAP mimicking the fBm $(Y_{t})$, c.f., Example 3.4.
Proposition 3.5.
Let $(S_{t}^{(n)})$ and $(A_{t}^{(n)})$ have mean functions $\mu_{S}$ and $\mu_{A}$, variance functions $\sigma_{S}^{2}$ and $\sigma_{A}^{2}$, and covariance functions $K_{S}$ and $K_{A}$, respectively. Then
$\mu_{I}(t):=\mathbb{E}[I_{t}^{(n)}]=\mu_{S}(t)+\mu_{A}(t),$ (3.5)
$\sigma_{I}^{2}(t):=\mathrm{Var}{[I_{t}^{(n)}]}=\sigma^{2}_{S}(t)+\sigma^{2}_{A}(t),$ (3.6)
$K_{I}(s,t):=\mathrm{Cov}[{I_{s}^{(n)},I_{t}^{(n)}}]=K_{S}(s,t)+K_{A}(s,t)=\sum\limits_{i=0}^{n}\sum\limits_{j=0}^{n}g_{i}(t)g_{j}(s)\mathrm{Cov}[X_{i},X_{j}]+K_{A}(s,t).$ (3.7)
The proof follows from the proof of Prop. 2.7. We introduce terminologies similar to the ones from the previous section.
Definition 3.6.
Let $(I_{t}^{(n)})$ be a RAP.
  1. 1.
    $(I_{t}^{(n)})$ is said to be a Gaussian RAP if its stochastic driver $(D_{t})$ is a Gaussian process.
  2. 2.
    $(I_{t}^{(n)})$ is said to be a standard RAP if its noise process $(A_{t}^{(n)})$ is a standard AP, and if for all $x\in[T_{0},T_{n}]$ and $j=1,\ldots,n$ it holds that $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$, and $g_{j}(x)\mathds{1}_{[T_{j-1},T_{j}]}(x)=f_{j}(x)\mathds{1}_{[T_{j-1},T_{j}]}(x)$.
Aside of simple cases, an $X$-RAP is not Markovian since no constraints are applied to the joint distribution of $X$. Hence, we introduce a relaxed notion of Markov property better suited to RAPs.
Definition 3.7.
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval, and $\tau_{0}<\tau_{1}<\ldots<\infty$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite, or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called $\tau$-conditionally-Markov if
$\mathbb{P}\left[Y_{t}\in\cdot\mid\mathcal{F}^{Y}_{s}\right]=\mathbb{P}\left[Y_{t}\in\cdot\mid Y_{\tau_{0}},\ldots,Y_{\tau_{m(s)}},Y_{s}\right]$ (3.8)
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, and $\tau_{m(s)}=\max\limits_{i\in\mathbb{N}}\{\tau_{i}\,|\,\tau_{i}\leqslant s\}$.
Now, we give a result on the conditional-Markov property of Gaussian RAPs similar to the one on the Markov property of Gaussian APs.
Theorem 3.8.
Let $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}=(S_{t}^{(n)}+A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ be a Gaussian $X$-RAP on $\{T_{0},T_{n}\}_{*}$. Then $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov if the following conditions are all satisfied:
  1. 1.
    The AP $(A_{t}^{(n)})$ is Markov, i.e., $K_{A}(s,t)=\sum\limits_{i=0}^{n-1}A_{1}(\min(s,t))A_{2}(\max(s,t))\mathds{1}_{(T_{i},T_{i+1})}(s,t)$.
  2. 2.
    For all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$,
    $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0,$ (3.9)
    $g_{j}(x)A_{1}(T_{j})\mathds{1}_{[T_{j-1},T_{j}]}(x)=A_{1}(x)\mathds{1}_{[T_{j-1},T_{j}]}(x).$ (3.10)
Proof.
Let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(I_{t}^{(n)})$ is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov if and only if
$\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m(s_{k})},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right]=\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m(s_{k})},I_{s_{k}}^{(n)}\right].$ (3.11)
where $m(s_{k}):=\max\limits\{i\in\mathbb{N}\,|\,T_{i}\leqslant s_{k}\}$. In the following, we will refer to $m(s_{k})$ by $m$ since $s_{k}$ is fixed.
We first show that $s_{1},\ldots,s_{k}$ can be selected to all be in the sub-interval $(T_{m},T_{m+1})$. To see this, assume there is an integer $j\in\{1,\ldots,k\}$ such that $s_{j}<T_{m}$ and $T_{m}<s_{j+1}$. Then
$\sigma(X_{0},\ldots,X_{m},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)})=\sigma(X_{0},\ldots,X_{m},A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)},I_{s_{j+1}}^{(n)},\ldots,I_{s_{k}}^{(n)})$ (3.12)
by Eq. (3.9). We also know that $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(X_{0},\ldots,X_{m})$ by the definition of the $X$-RAP, and $(A_{s_{1}}^{(n)},\ldots,A_{s_{j}}^{(n)})\perp\!\!\!\perp(I_{s_{j+1}}^{(n)},\ldots,I_{s_{k}}^{(n)},I_{t}^{(n)})$ since $(A_{t}^{(n)})$ is Markov. We conclude that
$\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right]=\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},I_{s_{j+1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right],$ (3.13)
which means we can assume that $s_{1},\ldots,s_{k}$ are all in the same sub-interval $(T_{m},T_{m+1})$.
Let us define $a_{m}(\cdot):=\sum\limits_{i=0}^{m}g_{i}(\cdot)X_{i}$, and
$\Delta_{q}:=\sum_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum_{i=1}^{k}c_{i,q}g_{m+1}(s_{i})X_{m+1}+\sum_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)},\quad q=1,\ldots,k-1,$ (3.14)
where the coefficients $(c_{i,q})$ are chosen such that
$\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0\quad\mbox{ and }\quad\mathbb{\det}\begin{pmatrix}c_{1,1}&\ldots&c_{1,k-1}\\ \vdots&&\vdots\\ c_{k-1,1}&\ldots&c_{k-1,k-1}\\ \end{pmatrix}\neq 0.$ (3.15)
This guarantees the following, where the notation ”$\,|\,(X_{0},\ldots,X_{m})$” means conditional on $(X_{0},\ldots,X_{m})$).
  1. 1.
    $\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},I_{s_{1}}^{(n)},\ldots,I_{s_{k}}^{(n)}\right]=\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m},\Delta_{1},\ldots,\Delta_{k-1},I_{s_{k}}^{(n)}\right]$.
  2. 2.
    $(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$ is a Gaussian vector. To see this, we observe that for all $q=1,\ldots k-1$,
    $\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0\implies\sum\limits_{i=1}^{k}c_{i,q}g_{m+1}(s_{i})=0,$ (3.16)
    where we used Eq. (3.10). Hence, $\Delta_{q}=\sum\limits_{i=1}^{k}c_{i,q}I_{s_{i}}^{(n)}=\sum\limits_{i=1}^{k}c_{i,q}a_{m}(s_{i})+\sum\limits_{i=1}^{k}c_{i,q}A_{s_{i}}^{(n)}$ for all $q=1,\ldots,k-1$, which implies that $(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$ is a Gaussian vector.
  3. 3.
    $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$, since $\sum\limits_{i=1}^{k}c_{i,q}K_{A}(s_{i},t)=0$.
  4. 4.
    $A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})$, since $(A_{t}^{(n)})$ is Markov.
To conclude, we need to show
$I_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})\text{ and }I_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m}).$ (3.17)
Since $(I_{t}^{(n)})=(S_{t}^{(n)}+A_{t}^{(n)})$, and $(S_{t}^{(n)})\perp\!\!\!\perp(A_{t}^{(n)})$, we have 3.17 if
$A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\,|\,(X_{0},\ldots,X_{m}),$ (3.18)
which is guaranteed by conditions 3.15. ∎
Remark 3.9.
If $(A_{t}^{(n)})$ is standard, see Definition 2.16, then Eq. (3.10) is equivalent to
$g_{j}(x)\mathds{1}_{[T_{j-1},T_{j}]}(x)=f_{j}(x)\mathds{1}_{[T_{j-1},T_{j}]}(x).$ (3.19)
This makes standard RAPs automatically conditionally-Markov.
Remark 3.10.
Depending on the coupling of $X$, the probability $\mathbb{P}[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m(s)},I_{s}^{(n)}]$ might simplify further. For instance, if $X$ has continuous marginals and is distributed according to Kantorovich’s coupling, then
$\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},\ldots,X_{m(s)},I_{s}^{(n)}\right]=\mathbb{P}\left[I_{t}^{(n)}\in\cdot\mid X_{0},I_{s}^{(n)}\right]$ (3.20)
because $X_{1},\ldots,X_{m(s)}$ are all deterministic functions of $X_{0}$.
Remark 3.11.
It is important to note that the conditional-Markov property is not symmetric in time. Define $\mathcal{G}_{t}^{I}=\sigma(I_{u}^{(n)}\,|\,t\leqslant u\leqslant T_{n})$, and let $s<t$ in $[T_{0},T_{n}]$. Then, to obtain
$\mathbb{P}\left[I_{s}^{(n)}\in\cdot\mid\mathcal{G}^{I}_{t}\right]=\mathbb{P}\left[I_{s}^{(n)}\in\cdot\mid I_{t}^{(n)},X_{k(t)},X_{k(t)+1},\ldots,X_{n}\right]$ (3.21)
where $k(t)=\min\limits\{i\in\mathbb{N}\,|\,T_{i}\geqslant t\}$, Condition (3.9) needs to be replaced by
$g_{j}(x)\mathds{1}_{[T_{j+1},T_{n}]}(x)=0,$ (3.22)
and Condition (3.10) by
$g_{j}(x)A_{1}(T_{j})\mathds{1}_{[T_{j},T_{j+1}]}(x)=A_{1}(x)\mathds{1}_{[T_{j},T_{j+1}]}(x)$ (3.23)
for all $j=0,\ldots,n-1$, and for all $x\in[T_{0},T_{n}]$.
Example 3.12.
We give an example of a non-standard $X$-RAP on $[T_{0},T_{2}]$ that is $\{T_{0},T_{2}\}_{*}$-conditionally-Markov, where $X_{0}\overset{\mathcal{L}}{=}X_{1}\overset{\mathcal{L}}{=}X_{2}\overset{\mathcal{L}}{=}\textrm{Uni}\,(\{-1,1\})$ are mutually independent. Consider the interpolating coefficients
$f_{0}(t)$ $=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{\left[T_{0},\frac{T_{1}+T_{2}}{2}\right]}(t)-\frac{T_{2}-t}{T_{1}-T_{0}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (3.24)
$f_{1}(t)$ $=\frac{t-T_{0}}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t)+\frac{T_{2}-t}{T_{2}-T_{1}}\mathds{1}_{(T_{1},T_{2}]}(t),$ (3.25)
$f_{2}(t)$ $=\frac{t-T_{1}}{T_{2}-T_{1}}\mathds{1}_{\left[T_{1},T_{2}\right]}(t).$ (3.26)
Let $(A_{t}^{(2)})_{t\in[T_{0},T_{2}]}$ be the arcade process built with these interpolating coefficients and driven by a standard Brownian motion. As shown in Example 2.18, $(A_{t}^{(2)})$ is Markov. For Eq. (3.9) to be satisfied, one only needs to impose $g_{2}(t)\mathds{1}_{[T_{0},T_{1}]}(t)=0$. For Eq. (3.10) to be satisfied, one requires that
$g_{1}(t)\mathds{1}_{[T_{0},T_{1}]}(t)=\frac{t-T_{0}}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t),$ (3.27)
$g_{2}(t)\mathds{1}_{\left[T_{1},\frac{T_{1}+T_{2}}{2}\right]}(t)=\left(\frac{t-T_{1}}{T_{2}-T_{1}}+\frac{(t-T_{1})T_{0}}{(T_{1}-T_{0})^{2}}\right)\mathds{1}_{\left[T_{1},\frac{T_{1}+T_{2}}{2}\right]}(t),$ (3.28)
$g_{2}(t)\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t)=\left(\frac{t-T_{1}}{T_{2}-T_{1}}+\frac{(T_{2}-t)T_{0}}{(T_{1}-T_{0})^{2}}\right)\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t).$ (3.29)
Outside of the considered intervals, the functions $g_{i}$ may take any values as long as they remain interpolating coefficients. Theorem 3.8 does not impose a condition on $g_{0}$. For example, we could choose $g_{i}=f_{i}$ outside the above intervals. Hence, all three conditions are fulfilled and this $X$-RAP is $\{T_{0},T_{2}\}_{*}$-conditionally-Markov. As we can see from the paths-simulation below, this process is different from a randomized stitched Brownian arcade on the second arc. The noise has been diminished to make the paths more informative. Simulating the signal function highlights the following: $X_{0}$ will determine the fate of the signal function on $[T_{1},T_{2}]$ since this RAP remembers the previously matched random variables when changing arc. On the first arc, where the process is simply a randomized Brownian bridge, to go from $X_{0}=-1$ to $X_{1}=-1$, for instance, there is only one way—a straight line. On the second arc, to go from $X_{1}=-1$ to $X_{2}=-1$, there are two ways. The evolution of the signal function is determined by the value that $X_{0}$ takes. This is illustrated by the paths of the signal function below: the blue path and the green path both take value $-1$ at $T_{1}$ and value $1$ at $T_{2}$, but have different values in $T_{0}$. Hence they differ on $[T_{1},T_{2}]$, as observed.
Figure 10: Paths simulation of the signal function of a non-standard $X$-RAP, where $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$ , c.f., Example 3.12.
Figure 11: Paths simulation of a non-standard $X$-RAP, where the noise is rescaled by 0.3 and $\{T_{0},T_{2}\}_{*}=\{0,5,10\}$, c.f., Example 3.12.
4 Filtered arcade martingales
In this section, given a martingale vector $X=(X_{0},\ldots,X_{n})$, i.e., a finite discrete-time martingale, we construct continuous-time martingales with respect to the filtration generated by an $X$-RAP, which interpolate between the convexly ordered components of the random vector $X$. These martingales solve an underlying stochastic filtering problem, and extend the martingale class of the information-based theory developed in [7], see also [8]. We call such martingales filtered arcade martingales (FAMs). We introduce the following notation:
  1. 1.
    $\mathcal{P}(\mathbb{R}^{n})$ is the set of Borel probability measures on $\mathbb{R}^{n}$, for $n\in\mathbb{N}_{0}$.
  2. 2.
    $\mathcal{P}_{1}(\mathbb{R}^{n})$ is the set of Borel probability measures on $\mathbb{R}^{n}$, for $n\in\mathbb{N}_{0}$, with finite first moment.
  3. 3.
[…]
    i
    =1
    ,,n+1
[…]
    1
[…]
    is the set of couplings of $\{\mu_{0},\mu_{1},\ldots,\mu_{n}\}\subseteq\mathcal{P}(\mathbb{R})$.
  4. 4.
    $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n}):=\{\pi\in\Pi(\mu_{0},\mu_{1},\ldots,\mu_{n})\cap\mathcal{P}_{1}(\mathbb{R}^{n+1})\,|\,(X_{0},X_{1},\ldots,X_{n})\sim\pi\implies\\ \mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m}]=X_{m}\text{ for all }m=0,\ldots,n\}$ is the set of martingale couplings of $\{\mu_{0},\mu_{1},\ldots,\mu_{n}\}\subseteq\mathcal{P}(\mathbb{R})$.
We recall that $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})\neq\emptyset$ if and only if $\int_{\mathbb{R}}x\,\mathrm{d}\mu_{0}(x)=\int_{\mathbb{R}}x\,\mathrm{d}\mu_{1}(x)=\ldots=\int_{\mathbb{R}}x\,\mathrm{d}\mu_{n}(x)$ and $\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{0}(x)\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{1}(x)\leqslant\ldots\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{n}(x)$ for any convex function $f$ on $\mathbb{R}$, i.e, the measures $\mu_{0},\mu_{1},\ldots,\mu_{n}$ are convexly ordered. In that case, we write $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}\leqslant_{\mathrm{cx}}\ldots\leqslant_{\mathrm{cx}}\mu_{n}$.
4.1 The one-arc FAM
Let $X=(X_{0},X_{1})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$, where $(\mu_{0},\mu_{1})\in\mathcal{P}_{1}(\mathbb{R})\times\mathcal{P}_{1}(\mathbb{R})$ and $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}$. It follows that $X_{0}\sim\mu_{0}$, $X_{1}\sim\mu_{1}$, and $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. Given an $X$-RAP $(I_{t}^{(1)})_{t\in[T_{0},T_{1}]}$ on the partition $\{T_{0},T_{1}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{1}]}$ such that $M_{T_{0}}\overset{a.s.}{=}X_{0}$ and $M_{T_{1}}\overset{a.s.}{=}X_{1}$. The BHM framework developed in [7] is recovered when $X_{0}=\mathbb{E}[X_{1}]=0$ and $T_{0}=0$.
Definition 4.1.
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FAM for $X\sim\pi^{X}$ on $[T_{0},T_{1}]$ is a stochastic process defined $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$.
The choice of the coupling $\pi^{X}$, alongside the interpolating coefficients and the driver of $(I_{t}^{(1)})$, dictate the behaviour of the paths of a FAM, since they tell us how $X_{0}$, $X_{1}$, and the information process $(I_{t}^{(1)})$ are related to each other.
Proposition 4.2.
The FAM $(M_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-martingale that interpolates between $X_{0}$ and $X_{1}$, almost surely.
Proof.
The process $(M_{t})$ is adapted to $(\mathcal{F}_{t}^{I})$ by construction. We show that it is a martingale that interpolates between $X_{0}$ and $X_{1}$ as follows.
  1. 1.
    $\mathbb{E}[|M_{t}|]<\infty$ for all $t\in[T_{0},T_{1}]$ by Jensen inequality, since $\mathbb{E}[|X_{1}|]<\infty$.
  2. 2.
    For $s<t$, $\mathbb{E}[M_{t}\,|\,\mathcal{F}_{s}^{I}]=M_{s}$ by the tower property of the conditional expectation.
  3. 3.
    $M_{T_{0}}=\mathbb{E}[X_{1}\,|\,X_{0}]\overset{a.s.}{=}X_{0}$, since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$ and $I_{T_{0}}^{(1)}=X_{0}$.
  4. 4.
    $M_{T_{1}}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{T_{1}}^{I}]\overset{a.s.}{=}X_{1}$, since $\sigma(X_{1})\subset\mathcal{F}_{T_{1}}^{I}$ by construction of $(I_{t}^{(1)})$.
Hence, $(M_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale that interpolates almost surely between $X_{0}$ and $X_{1}$ over the interval $[T_{0},T_{1}]$. ∎
Remark 4.3.
The process $(M_{t})_{t\in[T_{0},T_{1}]}$ is also a martingale with respect to its natural filtration, denoted $(\mathcal{F}_{t}^{M})_{t\in[T_{0},T_{1}]}$. For $s<t$, we have
$\mathbb{E}[M_{t}\,|\,\mathcal{F}^{M}_{s}]=\mathbb{E}[\mathbb{E}[M_{t}\,|\,\mathcal{F}^{M}_{s}]\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[M_{t}\,|\,\mathcal{F}^{I}_{s}]\,|\,\mathcal{F}^{M}_{s}]=\mathbb{E}[M_{s}\,|\,\mathcal{F}^{M}_{s}]=M_{s}.$ (4.1)
A relationship that we shall use later, between the RAP $(I_{t}^{(1)})$, its noise process $(A_{t}^{(1)})$, and its associated FAM $(M_{t})$ is the following:
Proposition 4.4.
If $I_{t}^{(1)}=g_{0}(t)X_{0}+g_{1}(t)X_{1}+A_{t}^{(1)}$ and $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$, then
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]$ (4.2)
for any pair $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. Furthermore, if $(A_{t}^{(1)})$ is Gauss-Markov with $K_{A}(x,y)=A_{1}(\min(x,y))A_{2}(\max(x,y))$, then
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
$\quad+\frac{A_{2}(t)}{A_{2}(s)}I_{s}+\mu_{A}(t)+\frac{A_{2}(t)}{A_{2}(s)}\mu_{A}(s)$ (4.3)
for any $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$.
Proof.
Let $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. We notice that
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$
$=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}],$ (4.4)
since $\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,X_{0},X_{1},\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]$, where we used the fact that $(A_{t}^{(1)})\perp\!\!\!\perp(X_{0},X_{1})$ by Definition 3.1. If $(A_{t}^{(1)})$ is Gauss-Markov, then
$\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]=\mathbb{E}[A_{t}^{(1)}\,|\,A_{s}^{(1)}]=\mu_{A}(t)+\frac{K_{A}(s,t)}{\sigma_{A}^{2}(s)}\left(A_{s}^{(1)}-\mu_{A}(s)\right).$ (4.5)
Hence, by the linearity property of the conditional expectation, we have
[…]
Moreover,
$I_{t}^{(1)}=\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{t}]=\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{t}]+g_{0}(t)X_{0}+g_{1}(t)M_{t},$ (4.7)
which implies
$\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{t}]=I_{t}^{(1)}-g_{0}(t)X_{0}-g_{1}(t)M_{t}.$ (4.8)
Then, inserting Eq. (4.8) into Eq. (4.6), and recalling that $K_{A}(s,t)=A_{1}(s)A_{2}(t)$, gives
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
$\quad+\frac{A_{2}(t)}{A_{2}(s)}I_{s}+\mu_{A}(t)+\frac{A_{2}(t)}{A_{2}(s)}\mu_{A}(s).$ (4.9)
If $(I_{t}^{(1)})$ is $\{T_{0},T_{1}\}$-conditionally-Markov, then $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$. In this case, the dynamics of $(M_{t})$ can be obtained using Bayes’ rule and Itô’s formula under mild assumptions. In the remainder of this section, we assume that the driver of $(I_{t}^{(1)})$ has a density function for $t\in[T_{0},T_{1}]$.
Proposition 4.5.
Let $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ be a one-arc FAM restricted to $t\in(T_{0},T_{1})$. Then
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)},$ (4.10)
where $f^{I_{t}^{(1)}\,|\,X_{0},X_{1}}$ is the conditional density of $(I_{t}^{(1)})$ given $X_{0}$ and $X_{1}$, and $F^{X_{1}\,|\,X_{0}}$ is the conditional distribution function of $X_{1}$ given $X_{0}$. In particular,
  1. 1.
    If $(X_{0},X_{1})$ is a continuous random vector, then
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\,|\,X_{0}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\,|\,X_{0}}(y)\,\mathrm{d}y}$ (4.11)
  2. 2.
    If $(X_{0},X_{1})$ is a discrete random vector, then
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\,|\,X_{0}]}{\sum\limits_{y}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\,|\,X_{0}]}.$ (4.12)
Proof.
By the Bayes rule,
$\mathbb{P}[X_{1}\leqslant y\,|\,X_{0},z\leqslant I_{t}^{(1)}\leqslant z+\epsilon]$ $=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\,|\,X_{0}]}{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0}]}$
$=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\,|\,X_{0}]}{\int_{\mathbb{R}}\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}=y]\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}$
$=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon,X_{1}\leqslant y\,|\,X_{0}]}{\int_{\mathbb{R}}\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}=y]\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.13)
This means that, by taking the limit when $\epsilon\rightarrow 0$,
$F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)=\frac{\frac{\,\mathrm{d}}{\,\mathrm{d}z}\mathbb{P}[I_{t}^{(1)}\leqslant z\,|\,X_{0},X_{1}\leqslant y]F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)},$ (4.14)
which implies
$\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)=\frac{f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.15)
Inserting the expression for $\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)$ into $M_{t}$, we obtain
$M_{t}=\int_{\mathbb{R}}y\,\mathrm{d}F^{X_{1}\,|\,X_{0},I_{t}^{(1)}}(y)=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.16)
Example 4.6.
Let $X_{0}\sim\mathcal{U}([-1,1])$, and $X_{1}\sim\mathcal{U}([-2,2])$. These probability distributions are convexly ordered, i.e., $X_{0}\leqslant_{\mathrm{cx}}X_{1}$, hence there exists at least one martingale coupling for $(X_{0},X_{1})$. We choose the coupling defined by
$X_{1}\,|\,X_{0}=\left\{\begin{array}[]{ll}\frac{3}{2}X_{0}+\frac{1}{2}&\mbox{with probability }\frac{3}{4},\\ \\ -\frac{1}{2}X_{0}-\frac{3}{2}&\mbox{with probability }\frac{1}{4}.\end{array}\right.$ (4.17)
This is a martingale coupling since $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. In fact, it can be shown that this coupling is the solution to a martingale optimal transport problem, see [25]. For any $\{T_{0},T_{1}\}$-conditionally-Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]=\frac{(9X_{0}+3)f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})-(X_{0}+3)f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}{6f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})+2f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}.$ (4.18)
Example 4.7.
Let $X_{0}\sim\mathcal{N}(0,1)$ and $X_{1}\sim\mathcal{N}(0,2)$, where $X_{1}\,|\,X_{0}\sim\mathcal{N}(X_{0},1)$. For any $\{T_{0},T_{1}\}$-conditionally-Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}.$ (4.19)
To simplify the expressions, we introduce the following notation:
$u(t,z,X_{0},y):=f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z),\quad u_{t}(t,z,X_{0},y):=\frac{\partial}{\partial t}u(t,z,X_{0},y),$ (4.20)
$u_{z}(t,z,X_{0},y):=\frac{\partial}{\partial z}u(t,z,X_{0},y),\quad u_{zz}(t,z,X_{0},y):=\frac{\partial^{2}}{\partial z^{2}}u(t,z,X_{0},y),$ (4.21)
$K_{\boldsymbol{\cdot}}(t,z,X_{0}):=\int_{\mathbb{R}}u_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y),\quad V_{\boldsymbol{\cdot}}(t,z,X_{0}):=\int_{\mathbb{R}}yu_{\boldsymbol{\cdot}}(t,z,X_{0},y)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y).$ (4.22)
Thus, under the conditions of Proposition 4.5, we may write Eq. (4.10) as
$M_{t}=\frac{V(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}.$ (4.23)
Proposition 4.8.
Let $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ be a one-arc FAM. If the $X$-RAP $(I_{t}^{(1)})$ is a semimartingale such that $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)}))$ where $N\subset(T_{0},T_{1})$ contains finitely many elements, then
$\,\mathrm{d}M_{t}=$ $\frac{V_{t}(t,I_{t}^{(1)},X_{0})-M_{t}K_{t}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}\,\mathrm{d}t+\frac{V_{z}(t,I_{t}^{(1)},X_{0})-M_{t}K_{z}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}\,\mathrm{d}I^{(1)}_{t}$
$+\Bigg{(}\frac{M_{t}K_{z}^{2}(t,I_{t}^{(1)},X_{0})-K_{z}(t,I_{t}^{(1)},X_{0})V_{z}(t,I_{t}^{(1)},X_{0})}{K^{2}(t,I_{t}^{(1)},X_{0})}$
$\hskip 142.26378pt+\frac{V_{zz}(t,I_{t}^{(1)},X_{0})-M_{t}K_{zz}(t,I_{t}^{(1)},X_{0})}{2K(t,I_{t}^{(1)},X_{0})}\Bigg{)}\,\mathrm{d}[I^{(1)}]_{t}.$ (4.24)
for $t\in(T_{0},T_{1})$.
Proof.
This is verified by a straightforward application of Itô’s formula. ∎
Remark 4.9.
The Itô condition, i.e., $(t,x)\rightarrow\frac{V(t,x,X_{0})}{K(t,x,X_{0})}$ is $C^{2}(((T_{0},T_{1})\setminus N)\times Im(I^{(1)})))$, where $N\subset(T_{0},T_{1})$ contains finitely many elements, imposes implicit integrability conditions on $(X_{0},X_{1})$.
Since Gaussian RAPs have been studied in detail above, we can specialize Proposition 4.8 to this particular subclass.
Corollary 4.10.
Under the conditions in Proposition 4.8, if the conditional probability distribution of $I_{t}^{(1)}$ given $(X_{0},X_{1})$ is $\mathcal{N}(g_{0}(t)X_{0}+g_{1}(t)X_{1}+\mu_{A}(t),\sigma_{A}^{2}(t))$, we have
$\frac{V_{t}(t,I_{t}^{(1)},X_{0})-M_{t}K_{t}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=-\frac{(\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]-M_{t}^{3})g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}$
$\hskip 56.9055pt+\Bigg{(}\frac{M_{t}g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)}{\sigma_{A}(t)}$
$\hskip 199.16928pt-\frac{(X_{0}g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t))g_{1}(t)}{\sigma_{A}^{2}(t)}\Bigg{)}\times\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.25)
$\frac{V_{z}(t,I_{t}^{(1)},X_{0})-M_{t}K_{z}(t,I_{t}^{(1)},X_{0})}{K(t,I_{t}^{(1)},X_{0})}=\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.26)
$\frac{M_{t}K_{z}^{2}(t,I_{t}^{(1)},X_{0})-K_{z}(t,I_{t}^{(1)},X_{0})V_{z}(t,I_{t}^{(1)},X_{0})}{K^{2}(t,I_{t}^{(1)},X_{0})}=\frac{I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)-g_{1}(t)M_{t}}{\sigma_{A}^{4}(t)}g_{1}(t)$
×$\hskip 341.43306pt\times\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}],$ (4.27)
$\frac{V_{zz}(t,I_{t}^{(1)},X_{0})-M_{t}K_{zz}(t,I_{t}^{(1)},X_{0})}{2K(t,I_{t}^{(1)},X_{0})}=\frac{\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}]-M_{t}^{3}}{2\sigma_{A}^{4}(t)}g_{1}^{2}(t)$
$\hskip 170.71652pt-\frac{2(I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t))+g_{1}(t)M_{t}}{2\sigma_{A}^{4}(t)}g_{1}(t)\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}].$ (4.28)
Proof.
Let $Z_{t}:=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}:=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$. The following computations of the partial derivatives $V_{t},K_{t},V_{z},V_{zz},K_{zz}$ in the case where $u(t,z,X_{0},y)=(1/\sqrt{2\pi\sigma_{A}^{2}(t)})\\ \exp[-(z-g_{0}(t)X_{0}-g_{1}(t)y-\mu_{A}(t))^{2}/(2\sigma^{2}_{A}(t))]$ lead to the result:
  1. 1.
    $V_{t}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]$
    $\quad+\frac{Z_{t}J_{t}\sigma_{A}(t)+\sigma_{A}^{\prime}(t)(Z_{t}^{2}-\sigma_{A}^{2}(t))}{\sigma_{A}^{3}(t)}V(t,I_{t}^{(1)},X_{0})$
    $\quad-\frac{g_{1}(t)(J_{t}\sigma_{A}(t)+2Z_{t}\sigma_{A}^{\prime}(t))-Z_{t}\sigma_{A}(t)g_{1}^{\prime}(t)}{\sigma_{A}^{3}(t)}K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}],$ (4.29)
  2. 2.
    $K_{t}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}(t)\sigma_{A}^{\prime}(t)-g_{1}^{\prime}(t)\sigma_{A}(t)}{\sigma_{A}^{3}(t)}g_{1}(t)K(t,I_{t}^{(1)},X_{0})\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]$
    $\quad+\frac{Z_{t}J_{t}\sigma_{A}(t)+\sigma_{A}^{\prime}(t)(Z_{t}^{2}-\sigma_{A}^{2}(t))}{\sigma_{A}^{3}(t)}K(t,I_{t}^{(1)},X_{0})$
    $\quad-\frac{g_{1}(t)(J_{t}\sigma_{A}(t)+2Z_{t}\sigma_{A}^{\prime}(t))-Z_{t}\sigma_{A}(t)g_{1}^{\prime}(t)}{\sigma_{A}^{3}(t)}K(t,I_{t}^{(1)},X_{0})M_{t},$ (4.30)
  3. 3.
    $V_{z}(t,I_{t}^{(1)},X_{0})=\frac{g_{1}(t)\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-Z_{t}V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{2}(t)},$ (4.31)
  4. 4.
    $K_{z}(t,I_{t}^{(1)},X_{0})=\frac{g_{1}(t)V(t,I_{t}^{(1)},X_{0})-Z_{t}K(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{2}(t)},$ (4.32)
  5. 5.
    $V_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{3}\,|\,X_{0},I_{t}^{(1)}]-2Z_{t}g_{1}(t)\mathbb{E}[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0})$
    $\quad+\frac{Z_{t}^{2}-\sigma_{A}^{2}(t)}{\sigma_{A}^{4}(t)}V(t,I_{t}^{(1)},X_{0}),$ (4.33)
  6. 6.
    $K_{zz}(t,I_{t}^{(1)},X_{0})$ $=\frac{g_{1}^{2}(t)E[X_{1}^{2}\,|\,X_{0},I_{t}^{(1)}]K(t,I_{t}^{(1)},X_{0})-2Z_{t}g_{1}(t)V(t,I_{t}^{(1)},X_{0})}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0})$
    $\quad+\frac{Z_{t}^{2}-\sigma_{A}^{2}(t)}{\sigma_{A}^{4}(t)}K(t,I_{t}^{(1)},X_{0}).$ (4.34)
Keeping the notation $Z_{t}=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}=X_{0}\,g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$ introduced in the proof, and introducing $U_{t}=\mathbb{E}[X_{1}^{3}\,|\,X_{0},I_{t}]-M_{t}^{3}$, the SDE satisfied by $(M_{t})$ can be rewritten as
$\,\mathrm{d}M_{t}=$ $U_{t}\left(\frac{g_{1}^{2}(t)}{2\sigma_{A}^{4}(t)}\,\mathrm{d}[I^{(1)}]_{t}-\frac{g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}\,\mathrm{d}t\right)+\frac{g_{1}(t)}{\sigma_{A}^{2}(t)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$
×$\times\Bigg{(}\Bigg{(}M_{t}\sigma_{A}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+Z_{t}\frac{\sigma_{A}(t)}{g_{1}(t)}\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)-J_{t}\Bigg{)}\,\mathrm{d}t$
$\hskip 142.26378pt-\frac{3g_{1}(t)M_{t}}{2\sigma_{A}^{2}(t)}\,\mathrm{d}[I^{(1)}]_{t}+\,\mathrm{d}I^{(1)}_{t}\Bigg{)}.$ (4.35)
If, furthermore, $(I_{t}^{(1)})$ is a standard RAP (see Definition 3.6), its driver $(D_{t})$ is Gauss-Markov with $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, where $H_{1}$ and $H_{2}$ are continuous functions on $[T_{0},T_{1}]$ such that $H_{1}/H_{2}$ is positive and non-decreasing on $[T_{0},T_{1}]$. Then, as shown in [14],
$[I^{(1)}]_{t}=[D]_{t}=\int_{T_{0}}^{t}H_{2}(s)\,\mathrm{d}H_{1}(s)-\int_{T_{0}}^{t}H_{1}(s)\,\mathrm{d}H_{2}(s),$ (4.36)
where the right-hand side is interpreted as a difference of Riemann-Stieltjes integrals. The right-hand side exists since
$\int_{T_{0}}^{t}H_{2}(s)\,\mathrm{d}H_{1}(s)-\int_{T_{0}}^{t}H_{1}(s)\,\mathrm{d}H_{2}(s)=\int_{T_{0}}^{t}H_{2}^{2}(s)\,\mathrm{d}\left(\frac{H_{1}}{H_{2}}\right)(s),$ (4.37)
and $H_{1}/H_{2}$ is continuous and monotone, so of bounded variation and differentiable almost everywhere.
Recalling that in the standard RAP case we have
$g_{1}(x)=\frac{H_{1}(x)H_{2}(T_{0})-H_{1}(T_{0})H_{2}(x)}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})},$
$A_{1}(x)=g_{1}(x)H_{2}(T_{1})$, and $A_{2}(x)=(H_{1}(T_{1})/H_{2}(T_{1}))H_{2}(x)-H_{1}(x)$, $\sigma_{A}^{2}(x)=A_{1}(x)A_{2}(x)$, we obtain the following expression for the SDE of the martingale $(M_{t})_{t\in[T_{0},T_{1}]}$.
Corollary 4.11.
Under the conditions of Proposition 4.8, if $(I_{t}^{(1)})$ is a standard $X$-RAP with driver covariance $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, then
$\,\mathrm{d}M_{t}=\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}$
×$\times\left(\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right),$ (4.38)
where $Z_{t}=I_{t}^{(1)}-g_{0}(t)X_{0}-\mu_{A}(t)$ and $J_{t}=X_{0}g_{0}^{\prime}(t)+\mu_{A}^{\prime}(t)$.
Proof.
The result follows from the following calculations:
$1.\quad\frac{g_{1}^{2}(t)(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{2\sigma_{A}^{4}(t)}-\frac{g_{1}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}}{\sigma_{A}(t)}=0,$ (4.39)
$2.\quad M_{t}\sigma_{A}(t)\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+Z_{t}\frac{\sigma_{A}(t)}{g_{1}(t)}\left(\left(\frac{g_{1}(t)}{\sigma_{A}(t)}\right)^{\prime}+g_{1}(t)\left(\frac{1}{\sigma_{A}(t)}\right)^{\prime}\right)$
$\hskip 227.62204pt-\frac{3g_{1}(t)M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{2\sigma_{A}^{2}(t)}$
$\quad\,\,\,=\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}.$ (4.40)
Introducing the notation
$\,\mathrm{d}N_{t}:=\left(\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}-J_{t}\right)\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t},$ (4.41)
we can then write
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}\,\mathrm{d}N_{t}.$ (4.42)
As shown below, the process $(N_{t})_{t\in[T_{0},T_{1}]}$, defined by the SDE 4.41 and initial condition $N_{T_{0}}=0$, is a martingale, and can be used to construct a standard Brownian motion on $[T_{0},T_{1}]$.
Definition 4.12.
Let $I\subseteq\mathbb{R}^{+}$ be an interval or the real positive line. A stochastic process $(Y_{t})_{t\in I}$ is said to be standard Brownian motion on an interval $[a,b]\subseteq I$ if there exists a standard Brownian motion $(B_{t})_{t\geqslant 0}$ such that $Y_{t}\overset{a.s.}{=}B_{t-a}$ for all $t\in[a,b]$.
Proposition 4.13.
Under the conditions in Proposition 4.8, the process $(W_{t})_{t\in[T_{0},T_{1}]}$ defined by
$W_{t}=\int_{T_{0}}^{t}\frac{1}{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\,\mathrm{d}N_{s}$ (4.43)
is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$.
Proof.
We begin with showing that $(N_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-martingale. We introduce the following notation: $h_{1}(t):=H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t)$, $h_{2}(t):=H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)$, $h_{3}(t):=H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})$, and
$S(X_{0},T_{0},t):=\int_{T_{0}}^{t}\frac{(g_{0}(u)X_{0}+\mu_{A}(u))h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$
Then, it follows that
$N_{t}=\int_{T_{0}}^{t}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u-S(X_{0},T_{0},t)-g_{0}(t)X_{0}-\mu_{A}(t)+I^{(1)}_{t}.$ (4.44)
Let $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s<t$. We next show that $\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]=N_{s}$. By the linearity property of the conditional expectation, we have
$\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]$ $=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u-M_{s}\int_{s}^{t}\frac{h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\quad-(S(X_{0},T_{0},s)+S(X_{0},s,t))-g_{0}(t)X_{0}-\mu_{A}(t)+\mathbb{E}[I^{(1)}_{t}\,|\,\mathcal{F}_{s}^{I}].$ (4.45)
By Proposition 4.4, the second and the last terms in the above expression of $\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]$ can be expressed as follows:
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
$\hskip 256.0748pt+\frac{A_{2}(t)}{A_{2}(s)}I_{s}^{(1)}+\mu_{A}(t)+\frac{A_{2}(t)}{A_{2}(s)}\mu_{A}(s),$ (4.46)
$\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u=X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\hskip 142.26378pt+\left(I_{s}^{(1)}-g_{1}(s)M_{s}-g_{0}(s)X_{0}+\mu_{A}(s)\right)\frac{1}{A_{2}(s)}\int_{s}^{t}\frac{A_{2}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$ (4.47)
Moreover, we notice that
$\frac{1}{A_{2}(s)}\int_{s}^{t}\frac{A_{2}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u=\frac{H_{1}(t)-H_{1}(s)-\frac{H_{1}(T_{1})}{H_{2}(T_{1})}\left(H_{2}(t)-H_{2}(s)\right)}{A_{2}(s)}=1-\frac{A_{2}(t)}{A_{2}(s)}.$ (4.48)
Hence,
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\quad\quad+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}+\mu_{A}(s)+\mu_{A}(t)$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+S(X_{0},s,t)+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}$
$\quad\quad+\mu_{A}(s)+\mu_{A}(t).$ (4.49)
Next, we observe that
$\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)-h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u$ $=\int_{s}^{t}\frac{H_{1}^{\prime}(u)H_{2}(T_{0})-H_{1}(T_{0})H_{2}^{\prime}(u)}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}\,\mathrm{d}u=g_{1}(t)-g_{1}(s),$ (4.50)
which allows one to write
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$=(g_{0}(t)-g_{0}(s))X_{0}+M_{s}\int_{s}^{t}\frac{h_{2}(u)}{h_{3}(u)}\,\mathrm{d}uM_{s}+S(X_{0},s,t)+I_{s}^{(1)}+\mu_{A}(s)+\mu_{A}(t).$ (4.51)
Inserting Eq. (4.1) into Eq. (4.1) gives:
$\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u-S(X_{0},T_{0},s)-g_{0}(s)X_{0}-\mu_{A}(s)+I^{(1)}_{s}=N_{s}.$ (4.52)
[…]
$[W]_{t}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[N]_{s}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[I^{(1)}]_{s}=\int_{T_{0}}^{t}\frac{h_{2}(s)}{h_{2}(s)}\,\mathrm{d}s=t-T_{0}.$ (4.53)
Since $(N_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale and $\mathbb{E}[[W]_{t}]<\infty$, $(W_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale. Hence, by Lévy’s characterization theorem, $(W_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$. ∎
The process $(W_{t})_{t\in[T_{0},T_{1}]}$ is referred to as the innovations process associated with the martingale $(M_{t})$. It follows that we can write $(M_{t})$ as an integral with respect its innovations process.
Corollary 4.14.
Under the conditions of Proposition 4.8, it holds that
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{s}^{(1)}]}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}.$ (4.54)
In the case that $(I^{(1)}_{t})$ is an $X$-randomized anticipative Brownian bridge on $[T_{0},T_{1}]$, that is,
$I_{t}^{(1)}=B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}(B_{T_{0}}-X_{0})-\frac{t-T_{0}}{T_{1}-T_{0}}(B_{T_{1}}-X_{1}),$ (4.55)
where $(B_{t})_{t\geq 0}$ is a standard Brownian motion, the expressions become significantly simpler.
Corollary 4.15.
If $(I^{(1)}_{t})_{t\in[T_{0},T_{1}]}$ is an $X$-randomized anticipative Brownian bridge on $[T_{0},T_{1}]$, the following holds:
  1. 1.
    The process $(M_{t})_{t\in[T_{0},T_{1}]}$ satisfies the SDE
    $M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{T_{1}-t}\left(\frac{I_{t}^{(1)}-M_{t}}{T_{1}-t}\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right).$ (4.56)
  2. 2.
    The process $(W_{t})_{t\in[T_{0},T_{1}]}$, defined by
    $W_{t}:=\int_{T_{0}}^{t}\frac{I_{u}^{(1)}-M_{u}}{T_{1}-u}\,\mathrm{d}u+I^{(1)}_{t}-X_{0},$ (4.57)
    is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$.
Proof.
It suffice to set $H_{1}(x)=x$, $H_{2}(x)=1$, $g_{0}(t)=(T_{1}-t)/(T_{1}-T_{0})$, and $g_{1}(t)=(t-T_{0})/(T_{1}-T_{0})$ in Eqs. (4.54) and (4.43). ∎
In case that $X_{0}=0$, $X_{1}$ is centered around 00, and $T_{0}=0$, the randomized anticipative Brownian bridge gives rise to the martingale developed in [7], i.e,
$M_{t}=\int_{0}^{t}\frac{\mathrm{Var}[X_{1}\,|\,I_{s}^{(1)}]}{T_{1}-s}\,\mathrm{d}W_{s}.$ (4.58)
Using Eq. (4.54), we give examples of FAMs where the underlying RAP is standard (see Definition 3.6).
Example 4.16.
Let $D_{t}=tB_{t}$ where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion. Then, $K_{D}(x,y)=\min(x,y)^{2}\max(x,y)$, and so $H_{1}(x)=x^{2}$, $H_{2}(x)=x$. The standard AP on $[T_{0},T_{1}]$ driven by $(D_{t})$ is given by
$A_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}D_{T_{0}}-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}D_{T_{1}}.$ (4.59)
We have
$K_{A}(x,y)=\frac{\min(x,y)(\min(x,y)-T_{0})\max(x,y)(T_{1}-\max(x,y))}{T_{1}-T_{0}},$ (4.60)
hence $A_{1}(x)=(x(x-T_{0}))/(T_{1}-T_{0})$ and $A_{2}(x)=x(T_{1}-x)$. A standard $X$-RAP with noise process $(A_{t}^{(1)})_{t\in[T_{0},T_{1}]}$ is
$I_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}(D_{T_{0}}-X_{0})-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}(D_{T_{1}}-X_{1}).$ (4.61)
We recall that a different interpolating coefficient for $X_{0}$ could have been chosen without disrupting the standard property of $(I_{t}^{(1)})$. The quadratic variation of $(I_{t}^{(1)})$ is given by $\,\mathrm{d}[I^{(1)}]_{t}=H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)\,\mathrm{d}t=t^{2}\,\mathrm{d}t$. Hence,
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]s}{T_{1}^{2}s-s^{2}T_{1}}\,\mathrm{d}W_{s}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{s}^{(1)}]}{T_{1}^{2}-sT_{1}}\,\mathrm{d}W_{s}.$ (4.62)
Example 4.17.
Let $(D_{t})$ be the Ornstein-Uhlenbeck process satisfying the SDE
$D_{t}=\int_{0}^{t}\theta\left(\mu-D_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\,\mathrm{d}B_{s},$ (4.63)
where $\theta>0,\sigma>0,\mu\in\mathbb{R}$. Then $K_{D}(s,t)=\frac{\sigma^{2}}{2\theta}\mathrm{e}^{\theta\min(s,t)}\mathrm{e}^{-\theta\max(s,t)}$, and so $H_{1}(x)=\\ \sigma^{2}/(2\theta)\exp{[\theta x]}$, $H_{2}(x)=\exp{[-\theta x]}$. The standard AP driven by $(D_{t})$ is
$A_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{0}}-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{1}}.$ (4.64)
The standard $X$-RAP with noise process $(A_{t}^{(1)})$ and the same interpolating coefficients for the signal function as in the noise process is
$I_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{0}}-X_{0})-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{1}}-X_{1}).$ (4.65)
The quadratic variation of $(I_{t}^{(1)})$ is given by $\,\mathrm{d}[I^{(1)}]_{t}=H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)\,\mathrm{d}t=\sigma^{2}\,\mathrm{d}t$. Hence, the associated FAM takes the form
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sigma\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{\frac{\sigma^{2}}{2\theta}\left(\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}\right)}\,\mathrm{d}W_{s}=X_{0}+\frac{2\theta}{\sigma}\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}}\,\mathrm{d}W_{s}.$ (4.66)
In general, one can expect that for a given coupling $\pi^{X}$, $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$ has no explicit analytical expression, even in the case where $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$. Nonetheless, we give an example where $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ and its SDE are explicit.
Example 4.18.
Let $X_{0}\sim\textrm{Uni}\,(\{-1,1\})$ and $X_{1}\sim\frac{1}{4}\delta_{-2}+\frac{1}{2}\delta_{0}+\frac{1}{4}\delta_{2}$ such that
$X_{1}\,|\,X_{0}=\left\{\begin{array}[]{ll}X_{0}+1&\mbox{with probability }\frac{1}{2},\\ \\ X_{0}-1&\mbox{with probability }\frac{1}{2}.\end{array}\right.$ (4.67)
Clearly, $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. Let $(B_{t})_{t\geqslant 0}$ be a standard Brownian motion and
$I_{t}^{(1)}=B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}(B_{T_{0}}-X_{0})-\frac{t-T_{0}}{T_{1}-T_{0}}(B_{T_{1}}-X_{1}).$ (4.68)
Moreover, let $\phi$ denote the density function of the measure $\mathcal{N}\left(0,\frac{(T_{1}-t)(t-T_{1})}{T_{1}-T_{0}}\right)$, and recall that $g_{1}(x)=(t-T_{0})/(T_{1}-T_{0})$. By Proposition 4.5, we have:
$M_{t}$ $=\frac{(X_{0}+1)\phi(I_{t}^{(1)}-X_{0}-g_{1}(t))\frac{1}{2}+(X_{0}-1)\phi(I_{t}^{(1)}-X_{0}+g_{1}(t))\frac{1}{2}}{\phi(I_{t}^{(1)}-X_{0}-g_{1}(t))\frac{1}{2}+\phi(I_{t}^{(1)}-X_{0}+g_{1}(t))\frac{1}{2}},$ (4.69)
$=X_{0}+\frac{\phi(I_{t}^{(1)}-X_{0}-g_{1}(t))-\phi(I_{t}^{(1)}-X_{0}+g_{1}(t))}{\phi(I_{t}^{(1)}-X_{0}-g_{1}(t))+\phi(I_{t}^{(1)}-X_{0}+g_{1}(t))},$ (4.70)
$=X_{0}+\tanh\left(\frac{I_{t}^{(1)}-X_{0}}{T_{1}-t}\right).$ (4.71)
The associated SDE is given by the integral form
[…]
Figure 12: Paths of the $X$-RAP $(I_{t}^{(1)})$, c.f., Example 4.18.
Figure 13: Paths of the associated FAM $(M_{t})$, c.f., Example 4.18.
4.2 Filtered arcade reverse martingales
Before moving on to the n-arc FAMs, one can apply the ideas developed so far to a different set of problems. Let $X=(X_{0},X_{1})$ be a random vector distributed according to a coupling $\pi^{X}\in\Pi(\mu_{0},\mu_{1})$, where $(\mu_{0},\mu_{1})\in\mathcal{P}_{1}(\mathbb{R})\times\mathcal{P}_{1}(\mathbb{R})$. Instead of predicting $X_{1}$ using the information revealed by an $X$-RAP, one might want to predict $X_{0}$ given different scenarios of the future generated by an $X$-RAP. We illustrate this idea with a brief thought experiment.
At the start of a farming season, a farmer might ask how the production of an agricultural produce should be kicked off so to achieve a target total production amount at the end of the season while minding the adverse weather forecast for the mid-season. If $X_{0}$ is the total amount produced by the end of the first day and $X_{1}$ is the total production at the end of the last day of the season, the farmer has a good idea about what $\pi^{X}$ is given his experience over many seasons. So, how should the farmer proceed? The answer is using a filtered arcade reverse martingale (FARM).
Definition 4.19.
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FARM for $X\sim\pi^{X}$ on $[T_{0},T_{1}]$ is a stochastic process defined by $M_{t}^{-}=\mathbb{E}[X_{0}\,|\,\mathcal{G}_{t}^{I}]$, where $\mathcal{G}_{t}^{I}:=\sigma(I_{u}^{(1)}\,|\,t\leqslant u\leqslant T_{1})$ and $t\in[T_{0},T_{1}]$.
As opposed to a FAM, a FARM is not a martingale, and it is not necessarily interpolating between $X_{0}$ and $X_{1}$, either. Instead, it is a reverse martingale (also called backward martingale) with respect to the reverse filtration $(\mathcal{G}_{t}^{I})$, which follows from the tower property of conditional expectation: $\mathbb{E}[M_{s}^{-}\,|\,\mathcal{G}_{t}^{I}]=\mathbb{E}[\mathbb{E}[X_{0}\,|\,\mathcal{G}_{s}^{I}]\,|\,\mathcal{G}_{t}^{I}]=\mathbb{E}[X_{0}\,|\,\mathcal{G}_{t}^{I}]=M_{t}^{-}$ for any $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. A FARM only interpolates between $X_{0}$ and $X_{1}$ if $(X_{0},X_{1})$ is a discrete-time reverse martingale, i.e., $\mathbb{E}[X_{0}\,|\,X_{1}]=X_{1}$, since $M_{T_{0}}^{-}=X_{0}$ and $M_{T_{1}}^{-}=\mathbb{E}[X_{0}\,|\,X_{1}]$ by construction.
To write a FARM $M_{t}^{-}$ as a function of $I_{t}^{(1)}$ and $X_{1}$, one needs another relaxation of the Markov property as hinted at in Remark 3.11.
Definition 4.20.
Let $\mathcal{I}\subseteq\mathbb{R}^{+}$ be a real interval, and $\tau_{0}>\tau_{1}>\ldots\geqslant\inf\mathcal{I}$ such that $\tau=\{\tau_{0},\tau_{1},\ldots\}\subset\mathcal{I}$. The set $\tau$ may be finite, or contain countably many elements. A stochastic process $(Y_{t})_{t\in\mathcal{I}}$ is called reverse $\tau$-conditionally Markov if
$\mathbb{P}\left[Y_{s}\in\cdot\mid\mathcal{G}^{Y}_{t}\right]=\mathbb{P}\left[Y_{s}\in\cdot\mid Y_{t},Y_{\tau_{m(t)}},\ldots,Y_{\tau_{0}}\right]$ (4.73)
for any $(s,t)\in\mathcal{I}^{2}$ such that $s\leqslant t$, where $\mathcal{G}^{Y}_{t}=\sigma(Y_{u}\,|\,t\leqslant u\leqslant\sup\mathcal{I})$ and $\tau_{m(s)}=\min\limits_{i\in\mathbb{N}}\{\tau_{i}\,|\,\tau_{i}\geqslant t\}$.
Similarly to the FAM, a FARM can be expressed by
$M_{t}^{-}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\,|\,X_{1}}(x)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{1},X_{0}=x}(I_{t}^{(1)})\,\mathrm{d}F^{X_{0}\,|\,X_{1}}(x)},$ (4.74)
for $t\in(T_{0},T_{1})$ if $(I_{t}^{(1)})$ is $\{T_{1},T_{0}\}$-reverse-Markov. In the previous section, we used standard RAPs to significantly simplify the SDEs that FAMs satisfy. For FARMs, the corresponding notion is the one of the reverse standard RAP.
Definition 4.21.
A RAP $(I_{t}^{(n)})$ is said to be a reverse standard RAP if its noise process $(A_{t}^{(n)})$ is a standard AP, and if for all $x\in[T_{0},T_{n}]$ and $j=0,\ldots,n-1$ it holds that $g_{j}(x)\mathds{1}_{[T_{j+1},T_{n}]}(x)=0$, and $g_{j}(x)\mathds{1}_{[T_{j},T_{j+1}]}(x)=f_{j}(x)\mathds{1}_{[T_{j},T_{j+1}]}(x)$.
Following the same procedure used for FAMs to derive their SDEs, while replacing the conditional-Markov property by the reverse conditional-Markov property, and standard RAPs by reverse standard RAPs, one finds that there exists a reverse standard Brownian motion $(W_{t}^{-})_{t\in[T_{0},T_{1}]}$ adapted to the reverse filtration $(\mathcal{G}_{t}^{I})$ such that
$M_{t}^{-}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{0}\,|\,I_{s}^{(1)},X_{1}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}^{-}.$ (4.75)
4.3 The n-arc FAM
Let $X=(X_{0},X_{1},\ldots,X_{n})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$, where $\mu_{i}\in\mathcal{P}_{1}(\mathbb{R})$ for $i=0,\ldots,n$ and $\mu_{i}\leqslant_{\mathrm{cx}}\mu_{i+1}$ for $i=0,\ldots,n-1$. Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ on the partition $\{T_{0},T_{n}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{n}]}$ such that $M_{T_{i}}\overset{a.s.}{=}X_{i}$, for $i=0,\ldots,n$. A major difference in the $n$-arc case is that $(I_{t}^{(n)})$ must be $\{T_{0},T_{n}\}_{*}$-conditionally-Markov to guarantee the interpolation of $(M_{t})$.
Definition 4.22.
Given an $X$-RAP $(I_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ that is $\{T_{0},T_{n}\}_{*}$-conditionally-Markov, an $n$-arc FAM for $X\sim\pi^{X}$ is a stochastic process of the form $M_{t}=\mathbb{E}[X_{n}\,|\,\mathcal{F}_{t}^{I}]=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,\\ X_{m(t)},I_{t}^{(n)}]$ where $m(t)=\max\{i\in\mathbb{N}\,|\,T_{i}\leqslant t\}$.
By the tower property of conditional expectation, $(M_{t})_{t\in[T_{0},T_{n}]}$ is a martingale with respect to $(\mathcal{F}_{t}^{I})_{t\in[T_{0},T_{n}]}$. Moreover, $M_{T_{i}}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{i}]=X_{i}$ since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})$. Hence, $(M_{t})$ is an interpolating martingale with respect to $(\mathcal{F}_{t}^{I})$. We notice that, without the conditional-Markov property, $M_{T_{i}}=\mathbb{E}[X_{n}\,|\,\mathcal{F}_{T_{i}}^{I}]$ is not necessarily a.s. equal to $X_{i}$. As for the case $n=1$, in the remainder of this section, we assume that the driver $(D_{t})$ of $(I_{t}^{(n)})$ has a density function on $[T_{0},T_{n}]$.
Proposition 4.23.
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM restricted to $(T_{0},T_{n})_{*}$. Then,
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\,\mathrm{d}F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}},$ (4.76)
where $F^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}$ is the distribution function of $X_{n}$ conditional on random vector $(X_{0},\ldots,\\ X_{m(t)})$. In particular:
  1. 1.
    If $(X_{0},\ldots,X_{m(t)},X_{n})$ is a continuous random variable, with density function $f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}$, then
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})f^{X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(y)\,\mathrm{d}y}.$ (4.77)
  2. 2.
    If $(X_{0},\ldots,X_{m(t)},X_{n})$ is a discrete random variable, then
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\,|\,X_{0},\ldots,X_{m(t)}]}{\sum\limits_{y}f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{n}=y}(I_{t}^{(n)})\mathbb{P}[X_{n}=y\,|\,X_{0},\ldots,X_{m(t)}]}.$ (4.78)
The proof follows along the lines of the one-arc case. Introducing the notation
$u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)=f^{I_{t}^{(n)}\,|\,X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1},X_{n}=y}(I_{t}^{(n)}),$ (4.79)
it is often more convenient to write
$M_{t}=\frac{\int_{\mathbb{R}^{n-m(t)}}y\,u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1},X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1},y)}{\int_{\mathbb{R}^{n-m(t)}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1},X_{n}\,|\,X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1},y)}.$ (4.80)
We can apply Itô’s formula in the same fashion as in the one-arc case. An interesting pattern appears if the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$, that significantly simplifies the $n$-arc case.
Proposition 4.24.
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM, where the signal function of $(I_{t}^{(n)})$ satisfies $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$. Then, $M_{t}=\mathbb{E}[X_{m(t)+1}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$.
Proof.
The result is straightforward for $t\in\{T_{0},T_{n}\}_{*}$, so we treat the case $t\in(T_{0},T_{n})_{*}$. Let $\varphi_{t}$ be the density function of $A_{t}^{(n)}$. Then,
$u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)$ $=\varphi_{t}\left(I_{t}^{(n)}-\sum_{i=0}^{m(t)}g_{i}(t)X_{i}-g_{m(t)+1}(t)x_{m+1})\right),$ (4.81)
since $g_{j}(x)\mathds{1}_{[T_{0},T_{j-1}]}(x)=0$ for all $j=1,\ldots,n$, and for all $x\in[T_{0},T_{n}]$. Denoting by $\tilde{F}$ the conditional distribution of $(X_{m(t)+1},\ldots,X_{n})$ given $(X_{0},\ldots,X_{m(t)})$, i.e.,
$\tilde{F}(x_{m+1},\ldots,x_{n})=F^{X_{m(t)+1},\ldots,X_{n}|X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n}),$ (4.82)
we have
$\int_{\mathbb{R}^{n-m(t)}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)$
$\hskip 85.35826pt=\int_{\mathbb{R}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1}|X_{0},\ldots,X_{m(t)}}(x_{m+1}).$ (4.83)
I also follows that
$\int_{\mathbb{R}^{n-m(t)}}y\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n})$
$=\int_{\mathbb{R}^{n-m(t)}}y\,\mathrm{d}F^{X_{n}|X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1}}(y)$
$\hskip 142.26378pt\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1}|X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1})$
$=\int_{\mathbb{R}^{n-m(t)-1}}\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},X_{m(t)+1}=x_{m+1},\ldots,X_{n-1}=x_{n-1}]$
$\hskip 142.26378pt\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1}|X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1})$
$=\int_{\mathbb{R}^{n-m(t)-1}}x_{n-1}\,\mathrm{d}F^{X_{m(t)+1},\ldots,X_{n-1}|X_{0},\ldots,X_{m(t)}}(x_{m+1},\ldots,x_{n-1}),$ (4.84)
where we use the martingale property of $(M_{t})$. Applying the same argument $m(t)$ times, one obtains
$\int_{\mathbb{R}^{n-m(t)}}y\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n})=\int_{\mathbb{R}}x_{m+1}\,\mathrm{d}F^{X_{m(t)+1}|X_{0},\ldots,X_{m(t)}}(x_{m+1}).$ (4.85)
Thus,
$\int_{\mathbb{R}^{n-m(t)}}yu(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)$
[…]
which leads to
$M_{t}$ $=\frac{\int_{\mathbb{R}^{n-m(t)}}yu(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)}{\int_{\mathbb{R}^{n-m(t)}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}\tilde{F}(x_{m+1},\ldots,x_{n-1},y)}$
$=\frac{\int_{\mathbb{R}}x_{m+1}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1}|X_{0},\ldots,X_{m(t)}}(x_{m+1})}{\int_{\mathbb{R}}u(I_{t}^{(n)},t,x_{m+1},\ldots,x_{n-1},y)\,\mathrm{d}F^{X_{m(t)+1}|X_{0},\ldots,X_{m(t)}}(x_{m+1})}$
$=\int_{\mathbb{R}}x_{m+1}\,\mathrm{d}F^{X_{m(t)+1}|X_{0},\ldots,X_{m(t)},I_{t}^{(n)}}(x_{m+1})$
$=\mathbb{E}[X_{m(t)+1}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}],$ (4.87)
where we use the Bayes rule in the second-last step. ∎
This allows one to use the results for the one-arc case, see Section 4.1 to derive the SDE for the $n$-arc FAM.
Proposition 4.25.
Let $M_{t}=\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}]$ be an $n$-arc FAM. If the $X$-RAP $(I_{t}^{(n)})$ is standard and a semimartingale, satisfying the condition in Proposition 4.24, with driver covariance $K_{D}(x,y)=H_{1}(\min(x,y))H_{2}(\max(x,y))$, such that $(t,x)\rightarrow\mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m(t)},I_{t}^{(n)}=x]$ is $C^{2}(((T_{0},T_{n})_{*}\setminus N)\times Im(I_{t}^{(n)}))$ where $N\subset(T_{0},T_{n})_{*}$ contains finitely many elements, then the following holds:
  1. 1.
    The process $(M_{t})_{t\in[T_{0},T_{n}]}$ satisfies the equation
    $M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s},$ (4.88)
    where
    $W_{t}=\int_{T_{0}}^{t}\frac{1}{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\,\mathrm{d}N_{s},$ (4.89)
    $\,\mathrm{d}N_{t}$ $=\bigg{(}\frac{Z_{t}(H_{1}^{\prime}(t)H_{2}(T_{m(t)+1})-H_{1}(T_{m(t)+1})H_{2}^{\prime}(t))-M_{t}(H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t))}{H_{1}(T_{m(t)+1})H_{2}(t)-H_{1}(t)H_{2}(T_{m(t)+1})}$
    $\hskip 284.52756pt-J_{t}\bigg{)}\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t},$ (4.90)
    and $Z_{t}=I_{t}^{(1)}-\sum\limits_{i=0}^{m(t)}g_{i}(t)X_{i}-\mu_{A}(t)$, and $J_{t}=\sum\limits_{i=0}^{m(t)}g_{i}^{\prime}(t)X_{i}+\mu_{A}^{\prime}(t)$.
  2. 2.
    The process $(W_{t})_{t\in[T_{0},T_{n}]}$ is an $(\mathcal{F}^{I}_{t})$-adapted standard Brownian motion on $[T_{0},T_{n}]$.
The proof follows from the one-arc case, see Proposition 4.8 to Corollary 4.14. Moreover, we note that
$M_{t}$ $=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}$
$=X_{m(t)}+\int_{T_{m(t)}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s))}\mathrm{Var}[X_{m(s)+1}\,|\,X_{0},\ldots,X_{m(s)},I_{t}^{(n)}]}{H_{1}(T_{m(s)+1})H_{2}(s)-H_{1}(s)H_{2}(T_{m(s)+1})}\,\mathrm{d}W_{s}.$ (4.91)
5 Outlook on IB-MOT and applications
In order to construct a FAM $(M_{t})$, one needs a martingale vector $X$ and an $X$-RAP. In some real-world situations, one might not have the vector $X$, but only its marginal distributions. In such a case, one must select a martingale coupling for these marginal distributions in order to construct a FAM. Could the behaviour of the FAMs associated with each possible coupling for the given marginal distributions inform us on which coupling to choose, i.e., is there an optimal coupling for a FAM? For each coupling, a FAM produces the best estimate of the last target random variable given the information generated by the $X$-RAP up until the time at which the estimation is performed. A natural question arises: Among all the estimates, which coupling yields the worst-case scenario? This idea of choosing a coupling based on paths connecting its marginal distributions on the space of probability measures reminds us of optimal transport. We flesh out our thinking by focusing on FAMs in the case of $n=1$.
Optimal transport (OT) dates back to Gaspard Monge in 1781 [23], with significant advancements by Leonid Kantorovich in 1942 [13] and Yann Brenier in 1987 [6]. It provides a way of comparing two measures, $\mu$ and $\nu$, defined on the Borel sets of topological Hausdorff spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We denote by $\mathcal{P}(\mathcal{X})$ and $\mathcal{P}(\mathcal{Y})$ the sets of Borel measures on $\mathcal{X}$ and $\mathcal{Y}$ respectively. A popular mental image in the context of optimal transport is a pile of sand, modelled by a measure $\mu$, and a hole, modelled by another measure $\nu$. One wishes to fill the hole with the sand in an optimal manner, by exercising the least amount of effort. To make this statement more precise, one needs a cost function $c:\mathcal{X}\times\mathcal{Y}\rightarrow[0,\infty]$ that measures the cost of transporting a unit mass from $x\in\mathcal{X}$ to $y\in\mathcal{Y}$. The optimal transport problem is concerned with how to transport $\mu$ to $\nu$ whilst minimizing the cost of transportation. That is, given $\mu\in\mathcal{P}(\mathcal{X})$ and $\nu\in\mathcal{P}(\mathcal{Y})$,
$\inf_{\pi\in\Pi(\mu,\nu)}\textup{{K}}(\pi):=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,\mathrm{d}\pi(x,y).$ (5.1)
This problem possesses many interesting properties. For instance, if $\mathcal{X}=\mathcal{Y}$ is a Polish space, and $(\mathcal{X},c)$ is a metric space, then
$W_{p}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{X}}c(x,y)^{p}\mathrm{d}\pi(x,y)\right)^{1/p}$ (5.2)
[…]
$\left(\frac{T_{1}-t}{T_{1}-T_{0}}X_{0}+\frac{t-T_{0}}{T_{1}-T_{0}}X_{1}\right)_{t\in[T_{0},T_{1}]}$ (5.3)
is the shortest paths from $\mu$ to $\nu$ in $(\mathcal{P}_{p}(\mathcal{X}),W_{p})$. This links optimal transport to interpolation on the space of random variables. There are two main differences between $(M_{t})$ and $(X_{0}(T_{1}-t)/(T_{1}-T_{0})+X_{1}(t-T_{0})/(T_{1}-T_{0}))$. The latter process is not a martingale, and does not incorporate noise, resulting in a deterministic interpolation conditionally on $(X_{0},X_{1})$. Could we modify OT in a way to incorporate noise, resulting in a martingale interpolator instead? We treat these two differences separately.
To achieve interpolation in the presence of noise within the optimal transport framework, the entropic regularization of optimal transport provides a way forward. Given $\mu\in\mathcal{P}(\mathcal{X})$, $\nu\in\mathcal{P}(\mathcal{Y})$, and $\varepsilon>0$, one considers
$\inf_{\pi_{\varepsilon}\in\Pi(\mu,\nu)}\textup{{K}}_{\varepsilon}(\pi_{\varepsilon}):=\inf_{\pi_{\varepsilon}\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\,\mathrm{d}\pi_{\varepsilon}(x,y)+\varepsilon\textup{ KL}(\pi_{\varepsilon}\,|\,\mu\otimes\nu)$ (5.4)
where KL is the Kullback-Leibler divergence
[…]
$F_{S}(\cdot)=\int_{\mathcal{X}}\int_{\mathcal{X}}F(x,y,\cdot)\,\mathrm{d}\pi_{S}^{*}(x,y),$ (5.6)
where $F(x,y,\cdot)$ is the distribution function of a Brownian bridge starting at time $T_{0}$ in value $x$ and ending at time $T_{1}$ in value $y$, is called a Schrödinger bridge. Such a process is the $X$-RAP
$I_{t}^{(1)}=B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}(B_{T_{0}}-X_{0})-\frac{t-T_{0}}{T_{1}-T_{0}}(B_{T_{1}}-X_{1})$ (5.7)
for $(X_{0},X_{1})\sim\pi_{S}^{*}$, and where $(B_{t})$ is a standard Brownian motion. Another characterization of such a process is the diffusion distributed according to $\mu$ at time $T_{0}$ and to $\nu$ at time $T_{1}$ with constant volatility $\sigma^{2}=\varepsilon/2(T_{1}-T_{0})$ and drift with the smallest possible $L^{2}(\Omega\times[T_{0},T_{1}])$-norm.
Next, we address how to create a martingale interpolator in the optimal transport setting. To this end, we adapt optimal transport to yield a martingale coupling: Given $\mu,\nu\in\mathcal{P}_{1}(\mathbb{R})$ in convex order,
$\inf_{\pi\in\mathcal{M}(\mu,\nu)}\textup{{K}}(\pi)=\inf_{\pi\in\mathcal{M}(\mu,\nu)}\int_{\mathbb{R}^{2}}c(x,y)\,\mathrm{d}\pi(x,y).$ (5.8)
We denote the minimizers of this problem by $\pi_{m}^{*}$. There are many differences between optimal transport and its martingale counterpart. For instance, the canonical cost function $c(x,y)=(x-y)^{2}$ for optimal transport cannot be used in the martingale context because, in this case, the objective function $\textup{{K}}(\pi)$ does not depend on $\pi$:
$\int_{\mathbb{R}^{2}}(x-y)^{2}\,\mathrm{d}\pi(x,y)$ $=\int_{\mathbb{R}^{2}}x^{2}+y^{2}-2xy\,\mathrm{d}\pi(x,y)$
$=\int_{\mathbb{R}^{2}}x^{2}\,\mathrm{d}\mu(x)+\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y)-2\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y)$
$=\int_{\mathbb{R}^{2}}x^{2}\,\mathrm{d}\mu(x)-\int_{\mathbb{R}^{2}}y^{2}\,\mathrm{d}\nu(y).$ (5.9)
Another difference is that one loses the geodesic interpretation since there is no counterpart to the Wasserstein distance in the martingale context.
Hence, when attempting to impose the martingale condition in OT, the interpolator corresponding to the coupling is lost, whereas when trying to inject noise in OT, the problem becomes the opposite of a martingale problem: one seeks to find a drift for a diffusion with a constant volatility. It is our view that FAMs could be used to reconcile the introduction of noise in martingale optimal transport (MOT) and martingale interpolation. Instead of comparing the distributions of $X_{0}$ and $X_{1}$ using the expectation $\mathbb{E}_{\pi}[c(X_{0},X_{1})]$, we propose to investigate the expectation of the cumulative cost in time on $[T_{0},T_{1}]$ incurred by estimating $X_{1}$ by $M_{t}$. This would result in an information-based version of MOT, since the information flow generated by the $X$-RAP is taken into account in determining the optimal coupling. Because we consider the cumulative cost in time on $[T_{0},T_{1}]$, we propose to multiply the cost by a weight function that informs the problem about the rate of convergence of $M_{t}$ towards $X_{1}$. By construction, the cost will always be small towards the end of the time interval, regardless of the coupling. The weight function allows for a fair comparison between the costs arising across different times by removing the bias introduced in the construction of $(M_{t})$.
This information-based version of MOT may be useful in an asset pricing context, for instance. A term structure of options allows to imply the market-consistent marginal distributions of the underlying asset at the fixed option expiry dates. Information-based MOT provides a method to explicitly construct the optimal no-arbitrage dynamics in continuous-time of an asset price process, e.g., the price process of a stock, that interpolates between the option-implied marginal distributions, i.e., the optimal FAM. By maximizing the expectation of the cumulative cost in time between $X_{1}$ and $M_{t}$ over the set of martingale couplings, and constructing the FAM corresponding to the maximizing martingale coupling, one obtains the ”worst case” peacock given the information generated by the underlying RAP, allowing us to establish probabilistic price bounds for the stock. The theory on, and applications thereof, information-based martingale optimal transport (IB-MOT) is developed in a follow-on paper [15].
Information-based stochastic interpolation can also be considered using FARMs $(M_{t}^{-})$. For example, a governmental body may need to decide on future climate goals in terms of, e.g., atmospheric CO2 emissions. Future goals depend on a succession of intermediate achievements, which are reached with uncertainty. So, these goals can be modelled by random variables with probability distributions in concave order, if a governmental body is interested in progressively decreasing the uncertainty around the mean until the final goal is reached. One can then look for an optimal reverse-martingale coupling linking the intermediate goals using the same procedure described above while replacing FAMs with FARMs: one aims at maximizing the expectation of the cumulative cost in time between $X_{0}$ and $M_{t}^{-}$ over the set of reverse-martingale couplings. Then, constructing the FARM corresponding to the maximizing reverse-martingale coupling, one obtains the ”worst case” reverse-peacock given the information generated by the underlying RAP, allowing one to understand if the goals are likely to be achieved, and if so, what trajectories must be taken in the near future to succeed.
6 Conclusions
Given a finite sequence of convexly ordered random variables that are indexed by pre-specified fixed dates, we show how continuous-time martingales, the so-called filtered arcade martingales (FAMs, Definitions 4.1 and 4.22) can be constructed explicitly, such that the target random variables are matched, at the given dates, almost surely. These interpolating martingales, defined as conditional expectations, filter noisy information about the last target random variable such that the interpolation through the whole sequence of random variables is guaranteed. This martingale-interpolating method relies on a specific class of stochastic processes, the randomized arcade processes (RAPs, Definition 3.1). These processes, which interpolate strongly between the target random variables, generate information flows that reveal each target random variable at its indexing date, thus allowing a FAM to interpolate through the sequence of random variables. FAMs are solutions to filtering problems where the best estimate of the last target random variable is calculated given the filtration generated by the observation process, i.e., the RAP.
A randomized arcade process is the sum of two independent components: a signal function and a noise process, the latter is the so-called arcade process (AP, Definition 2.2). An AP is a functional of a stochastic driving process, and interpolates strongly between zeros. A well-known example of an AP over one period of time is the anticipative Brownian bridge (Example 2.3). Since APs play the role of noise, a natural class of APs are the ones driven by a Gaussian process. We give sufficient condition for such an AP to be Markovian (Corollary 2.11), and show that there are Markovian APs that are not anticipative representations of Markov bridges (Example 2.18). As for the signal function, it interpolates between the target random variables, and it is deterministic conditional on these random variables. Since a RAP is a sum of a signal function and an AP, it does not possess the Markov property with respect to its natural filtration, apart from in trivial cases. Instead, it may have a similar property, the conditionally-Markov property (Definition 3.7), which plays a central role for the computational tractability of a FAM, and for guaranteeing the interpolation in the case with more than two random variables. We give sufficient conditions for a RAP to satisfy the conditionally-Markov property (Theorem 3.8), and we consider FAMs constructed with such RAPs. Using the Bayes rule and Itô’s lemma, we derive the SDEs satisfied by such FAMs (Propositions 4.5 and 4.8). If furthermore, the RAPs are standard (Definition 3.6), the obtained FAMs can be expressed in terms of an integral with respect to a Brownian motion adapted to the filtration generated by the RAP underlying the FAM (Proposition 4.13). This Brownian motion corresponds to the innovations process in stochastic filtering theory.
The FAM theory relies on convexly ordered probability measures, that is, the existence of a martingale coupling for the target variables. In some real life scenarios, one might not have the coupling, but only its marginal distributions. In that case, one must select a martingale coupling for these marginal distributions in order to construct a FAM. As a potential application in the field of optimal transport, that will be studied in a followup paper, we propose the use of the paths of FAMs to select an optimal coupling for the marginal distributions. This method is inspired by martingale optimal transport and Schrödinger’s problem. It aims at introducing noise in MOT in a similar fashion to how the entropic regularization of optimal transport introduces noise in optimal transport. The noise contained in FAMs, which depends on the choice of the underlying RAP, is taken into consideration in the selection of an optimal coupling for the target random variables: we aim at maximizing the expectation of the cumulative cost in time between the final target random variable and the considered FAM over the set of martingale couplings. The same ideas can be applied to FARMs, where we look instead for an optimal reverse-martingale coupling based on the paths of FARMs. Arcade processes and randomized arcade processes, alongside the filtered arcade martingales, appear thereof to suggest applications in, e.g., financial and insurance mathematics, mathematical biology, statistics, and climate science.
Acknowledgments
A. Macrina is grateful for discussions on stochastic interpolation with Camilo A. Garcia Trillos and Jun Sekine, especially in the context of information-based modelling and stochastic filtering, during a research visit at Osaka University in March 2017 as part of a Daiwa Anglo-Japanese Foundation Small Grant, Ref: 3078/12370. Osaka University’s hospitality is acknowledged and appreciated. A. Macrina is also grateful for the support by the Fields Institute for Research in Mathematical Sciences through the award of a Fields Research Fellowship in 2022. The contents of this manuscript are solely the responsibility of the authors and do not necessarily represent the official views of the Fields Institute. G. Kassis acknowledges the UCL Department of Mathematics for a Teaching Assistantship Award. Furthermore, the authors thank L. P. Hughston for helpful comments and suggestions at an early stage of the work on arcade processes, B. Acciaio and G. Pammer for pointers to optimal transport and stretched Brwonian motion, S. Cohen, J. Guyon, F. Krach, J. Obloj, G. W. Peters, and T.-K. L. Wong for useful conversations, and J. Akahori, K. Owari and D. C. Schwarz for feedback. Attendees of the AIFMRM Research Seminar at the University of Cape Town (Aug. 2021) and participants in the SIAM Conference on Mathematics of Data Science (MDS22, Sept. 2022), the 7th International Conference Mathematics in Finance—MiF Kruger Park— (July 2023, South Africa), the Talks in Financial and Insurance Mathematics at ETH Zurich (Nov. 2023, Switzerland), the Department of Mathematics Seminar of Ritsumeikan University (Nov. 2023, Japan), and participants in the FAMiLLY Workshop at the University of Liverpool (Dec. 2023, U.K.) and the XXV Workshop on Quantitative Finance (Apr. 2024, Italy) are thanked for their comments and suggestions.